### v3

# Upgraded free independence phenomena for random unitaries

Abstract.
We study upgraded free independence phenomena for unitary elements $u_{1}$, $u_{2}$, …representing the large-$n$ limit of Haar random unitaries, showing that free independence extends to several larger algebras containing $u_{j}$ in the ultraproduct of matrices $\prod_{n\to{\mathcal{U}}}M_{n}({\mathbb{C}})$. Using a uniform asymptotic freeness argument and volumetric analysis, we prove free independence of the Pinsker algebras ${\mathcal{P}}_{j}$ containing $u_{j}$. The Pinsker algebra ${\mathcal{P}}_{j}$ is the maximal subalgebra containing $u_{j}$ with vanishing $1$-bounded entropy [27]; ${\mathcal{P}}_{j}$ in particular contains the relative commutant $\{u_{j}\}^{\prime}\cap\prod_{n\to{\mathcal{U}}}M_{n}({\mathbb{C}})$, more generally any unitary that can be connected to $u_{j}$ by a sequence of commuting pairs of Haar unitaries, and any unitary $v$ such that $v{\mathcal{P}}_{j}v^{*}\cap{\mathcal{P}}_{j}$ is diffuse. Through an embedding argument, we go back and deduce analogous free independence results for ${\mathcal{M}}^{{\mathcal{U}}}$ when ${\mathcal{M}}$ is a free product of Connes embeddable tracial von Neumann algebras ${\mathcal{M}}_{i}$, which thus yields (in the Connes-embeddable case) a generalization and a new proof of Houdayer–Ioana’s results on free independence of approximate commutants [32]. It also yields a new proof of the general absorption results for Connes-embeddable free products obtained by the first author, Hayes, Nelson, and Sinclair [31].
1991 Mathematics Subject Classification:
46L54, 46L10, 60B20, 94A17
1. Introduction
1.1. Main results
A fundamental result in random matrix theory is that for independent $n\times n$ Haar random unitaries $U_{1}^{(n)}$, $U_{2}^{(n)}$, …, the trace of any word in the $U_{j}^{(n)}$’s and their adjoints converges to trace of the corresponding word in the free group [60, 63]. Thus, the von Neumann algebra of the free group $L(F_{\infty})$ arises from the large $n$ limit of random matrices. Recently, Houdayer and Ioana [32] discovered that free independence of algebras ${\mathcal{M}}_{i}$ in a free product ${\mathcal{M}}=*_{i\in I}{\mathcal{M}}_{i}$ can be upgraded to free independence of their commutants in the ultrapower of ${\mathcal{M}}$. Thus, for instance, if $b_{1}$, $b_{2}$, …are elements in $L(F_{\infty})$ such that $b_{j}$ approximately commutes with the group generator $g_{j}$, then $b_{1}$, $b_{2}$, …must be approximately freely independent. Thus, it is natural to ask whether an analogous statement holds for Haar random unitary matrices. We will give an affirmative answer to this question, so that matrix ultraproducts exhibit similar behavior to free products in this regard, and in addition prove several generalizations.
Theorem A (Asymptotic freeness of approximate commutants).
Let $U_{1}^{(n)}$, $U_{2}^{(n)}$, …be independent $n\times n$ Haar random unitary matrices. Let $B_{1}^{(n)}$, …, $B_{m}^{(n)}$ be random matrices on the same probability space such that $\lVert B_{j}^{(n)}\rVert\leq 1$ and $\lim_{n\to\infty}\lVert[U_{j}^{(n)},B_{j}^{(n)}]\rVert_{2}=0$ almost surely for $j=1,\dots,k$. Then $B_{1}^{(n)}$, …, $B_{k}^{(n)}$ are almost surely asymptotically freely independent.
Here also each matrix $B_{j}^{(n)}$ can be replaced by a tuple; see Theorem 2.12. The challenge of Theorem A is that $B_{j}^{(n)}$ is allowed to depend arbitrarily on the $U_{j}^{(n)}$. Thus, the proof requires a version of Voiculescu’s asymptotic freeness [63, Corollary 2.7] that applies uniformly to all the possible values of $B_{j}^{(n)}$. After considering the diagonalization of the $U_{j}^{(n)}$’s (see §2.1), we will show a uniform asymptotic freeness result for matrices $B_{j}^{(n)}$ that are asymptotically supported in $\epsilon n$-bands around the diagonal (Lemma 2.10). To guarantee each moment condition, we test it on a $\delta$-dense subset for some small $\delta$ by playing off the exponential concentration of measure for the Haar random unitary matrices against the small dimension of the $\epsilon$-bands compared to the ambient matrix space.
Such asymptotic results can be conveniently formulated using ultraproducts of tracial von Neumann algebras (see [8, Appendix A] for background); intuitively, elements of the ultraproduct $\prod_{n\to{\mathcal{U}}}{\mathcal{M}}_{n}$ capture all possible limiting behaviors of elements $x_{n}$ from ${\mathcal{M}}_{n}$, and thus allow asymptotic or approximate statements to be reformulated as exact statements. Letting ${\mathcal{U}}$ be a free ultrafilter on ${\mathbb{N}}$, Houdayer–Ioana’s result in the special case of $L(F_{\infty})$ would say that the commutants $\{g_{j}\}^{\prime}\cap L(F_{\infty})^{{\mathcal{U}}}$ are freely independent of each other. Meanwhile, the ultraproduct version of Theorem A is the following.
Theorem B (Freeness of relative commutants).
Let ${\mathcal{U}}$ be a free ultrafilter on ${\mathbb{N}}$, and let ${\mathcal{Q}}=\prod_{n\to{\mathcal{U}}}M_{n}({\mathbb{C}})$ be the ultraproduct of matrix algebras. Let $U_{1}^{(n)}$, $U_{2}^{(n)}$, …be independent $n\times n$ Haar random unitary matrices on a probability space $(\Omega,\mathcal{F},P)$. For each $\omega\in\Omega$, let $u_{j}(\omega)=[U_{j}^{(n)}(\omega)]_{n\in{\mathbb{N}}}\in\mathcal{Q}$ be the corresponding element of the matrix ultraproduct. Then almost surely
$\{u_{1}(\omega)\}^{\prime}\cap\mathcal{Q},\quad\{u_{2}(\omega)\}^{\prime}\cap{\mathcal{Q}},\quad\dots$
are freely independent.
This method of volumetric analysis on the space of matrices goes back at least to von Neumann [64]. Voiculescu formalized the exponential growth rate of volumes of the set of matrix microstate with certain moments through his free entropy $\chi$ [61], which he used to show that the free group von Neumann algebra has no Cartan subalgebra [62] (see also [21]). Jung [37] defined the related notion of strong $1$-boundedness, a condition of ‘‘small microstate dimension’’ which is independent of the choice of generators of the von Neumann algebra, and which Hayes later captured through the metric-entropy invariant $h$ [27]. Volumetric analysis and high-dimensional concentration of measure in random matrix theory [4, 39] form a powerful combination with many applications to the structure of $L(F_{\infty})$ [31, 28, 30]. These techniques also relate closely to deep questions about the growth rates of approximate representations [50, 52, 51].
Using the toolkit of $1$-bounded entropy and a small amount of model theory for von Neumann algebras [15, 16, 34], we show that this upgrading of freeness phenomenon applies not only to the commutants of $u_{j}$ in ${\mathcal{Q}}$, but to the much larger Pinsker algebra of $u_{j}$. A Pinsker algebra [31, Definition after Theorem B] in a von Neumann algebra ${\mathcal{M}}$ is a maximal von Neumann algebra ${\mathcal{P}}\subseteq{\mathcal{M}}$ such that $h({\mathcal{P}}:{\mathcal{M}})=0$ (this terminology is motivated by an analogous construction in ergodic theory). Thanks to [27, Lemma A.12], every diffuse von Neumann subalgebra ${\mathcal{A}}\subseteq{\mathcal{M}}$ with $h({\mathcal{A}}:{\mathcal{M}})=0$ is contained in a unique Pinsker algebra ${\mathcal{P}}$. Moreover, the general properties of $1$-bounded entropy (see e.g. [31, §1.2], [29, §2.3]) imply the following useful properties of the Pinsker algebra ${\mathcal{P}}$. If ${\mathcal{B}}\supseteq{\mathcal{A}}$ is amenable or has property Gamma, then ${\mathcal{B}}$ must be contained in ${\mathcal{P}}$. If $u$ is unitary and $u{\mathcal{P}}u^{*}\cap{\mathcal{P}}$ is diffuse, then $u$ must be contained in ${\mathcal{P}}$, and thus in particular ${\mathcal{P}}$ contains many different weakened versions of normalizers as in [33, 49, 14, 19]. In fact, $L^{2}({\mathcal{M}})\ominus L^{2}({\mathcal{P}})$ is a coarse bimodule over ${\mathcal{P}}$ [27, Theorem 3.8].
Theorem C (Freeness of Pinsker algebras).
Let $u_{j}(\omega)$ be as in Theorem B. Let ${\mathcal{P}}_{j}$ be the Pinsker algebra of $u_{j}(\omega)$. Then almost surely ${\mathcal{P}}_{1}$, ${\mathcal{P}}_{2}$, …are freely independent.
This immediately implies the following corollary, for instance.
Corollary D (Freeness of amenable algebras).
Let $u_{j}(\omega)$ be as in Theorem B. Then almost surely the following statement holds: If ${\mathcal{A}}_{j}$ is any amenable subalgebra containing $u_{j}(\omega)$, then ${\mathcal{A}}_{1}$, ${\mathcal{A}}_{2}$, …are freely independent.
Another consequence of Theorem C is freeness of the sequential commutation orbits of $u_{j}$ studied by [38]; see also [20], [10]. Recall a Haar unitary in a tracial von Neumann algebra ${\mathcal{M}}$ is any unitary element $u$ satisfying $\operatorname{tr}^{{\mathcal{M}}}(u^{m})=0$ for all $m\in{\mathbb{Z}}\setminus\{0\}$, or equivalently a unitary $u$ whose spectral measure is the Haar measure on the circle. [^1] This sense of Haar unitary is not to be confused with the Haar random unitary matrix $U^{(n)}$ which is a random matrix chosen according to the Haar measure on the $n\times n$ unitary group. For this reason, we will always refer to the latter as a Haar random unitary. We denote the set of Haar unitaries by $\mathcal{H}({\mathcal{M}})$. Following [38], for $u,v\in\mathcal{H}({\mathcal{M}})$, we say $u\sim_{k}v$ if there exist Haar unitaries $u=u_{0},u_{1},\dots,u_{k}=v$ in an ultrapower ${\mathcal{M}}^{{\mathcal{U}}}$ such that $[u_{j-1},u_{j}]=0$ for $j=1,\dots,k$. We also write $u\sim v$ if $u\sim_{k}v$ for some $k$.
The sequential commutation orbit of $u$ is its equivalence class under the relation $\sim$. As a consequence of [38, Fact 2.9], the sequential commutation orbit of a Haar unitary $u$ is always contained inside the Pinsker algebra of $u$. (Actually, in the special case of $L(F_{\infty})$, the Pinsker algebra of some ${\mathcal{A}}$ with $h({\mathcal{A}}:L(F_{\infty}))=0$ is equal to the algebra generated by its sequential commutation orbit [38, Theorem 5.4] as a consequence of the recent resolution of the Peterson-Thom conjecture of [49, §7.5], which occurred through a combination of $1$-bounded entropy techniques [28] and strong convergence of tensor product random matrix models [3, 7, 11, 46, 9].)
Corollary E (Freeness of sequential commutation orbits).
Let $u_{j}(\omega)\in\mathcal{Q}=\prod_{n\to{\mathcal{U}}}M_{n}(\mathbb{C})$ be as in the Theorem B. Almost surely, the von Neumann algebras generated by the sequential commutation orbits of $u_{1}(\omega)$, $u_{2}(\omega)$, …respectively are freely independent.
Although Theorem C focuses on the matrix ultraproduct, we are able to transfer these results to $L(F_{\infty})$ and more generally to arbitrary free products of Connes-embeddable von Neumann algebras through a natural argument using embeddings and countable saturation (a concept from model theory).
Theorem F (Freeness phenomena in ultrapowers of free products).
Let $({\mathcal{M}}_{i})_{i\in I}$, …be diffuse Connes-embeddable tracial von Neumann algebras, and let ${\mathcal{M}}=*_{i\in I}{\mathcal{M}}_{i}$. Let ${\mathcal{V}}$ be a free ultrafilter on some index set $J$. Then
  1. (1)
    If ${\mathcal{A}}_{i}\subseteq{\mathcal{M}}^{{\mathcal{V}}}$ with $h({\mathcal{A}}_{i}:{\mathcal{M}}^{{\mathcal{V}}})=0$ and ${\mathcal{A}}_{i}\cap{\mathcal{M}}_{i}$ diffuse for each $i$, then the algebras $({\mathcal{M}}_{i}\vee{\mathcal{A}}_{i})_{i\in I}$ are freely independent.
  2. (2)
    Let ${\mathcal{C}}_{i}\subseteq{\mathcal{M}}^{{\mathcal{V}}}$ be the algebra generated by the sequential commutation orbits of Haar unitaries in ${\mathcal{M}}_{i}$. Then $({\mathcal{C}}_{i})_{i\in I}$ are freely independent. In particular, if $u_{i}\in{\mathcal{M}}_{i}$ are diffuse unitaries, then $\{u_{i}^{\prime}\cap{\mathcal{M}}^{{\mathcal{V}}}\}_{i\in I}$ are freely independent.
  3. (3)
    Let ${\mathcal{N}}_{i}\subseteq{\mathcal{M}}^{{\mathcal{V}}}$ be the von Neumann algebra generated by the wq-normalizer of ${\mathcal{M}}_{i}$ in ${\mathcal{M}}^{{\mathcal{V}}}$ as defined in [19], namely,
    ${\mathcal{N}}_{i}=\mathrm{W}^{*}(\{u\in\mathcal{U}({\mathcal{M}}^{{\mathcal{V}}}):u{\mathcal{M}}_{i}u^{*}\cap{\mathcal{M}}_{i}\text{ diffuse}\}).$
    Then $({\mathcal{N}}_{i})_{i\in I}$ are freely independent.
Note that this theorem is not stated in terms of Pinsker algebras because we have no assumptions on the ${\mathcal{M}}_{i}$ other than Connes embeddability, so we have not assumed that $h({\mathcal{M}}_{i})=0$. In fact, in the proof, we embed ${\mathcal{M}}_{i}$ into ${\mathcal{R}}^{{\mathcal{U}}}$ and study Pinsker algebras of these copies of ${\mathcal{R}}^{{\mathcal{U}}}$ inside their free product. One can deduce that the algebras ${\mathcal{A}}_{i}=\bigvee_{u\in\mathcal{H}({\mathcal{M}}_{i})}{\mathcal{P}}_{u}$ generated by the Pinsker algebras of Haar unitaries in ${\mathcal{M}}_{i}$ are freely independent. Alternatively, it would be natural to state the theorem in terms of Pinsker algebras relative to ${\mathcal{M}}_{i}$, that is, maximal algebras with entropy 00 conditioned on ${\mathcal{M}}_{i}$ (see [34, §6]), though for the sake of length we do not develop this perspective fully here.
Theorem F in particular gives a new proof and generalization of the freeness result for commutants of Houdayer-Ioana [32, Theorem B] in the case of Connes embeddable free products. However, we cannot handle the case of amalgamated free products through this method. Indeed, it is not even known if free products with amalgamation over a non-amenable algebra preserve Connes embeddability, and even if the amalgam is diffuse amenable, we expect similar issues to arise as in [31, Remark 5.13].
Theorem F (1) also recovers [31, Theorem A], which says that if ${\mathcal{A}}_{i}$ is contained in ${\mathcal{M}}$ with ${\mathcal{A}}_{i}\cap{\mathcal{M}}_{i}$ diffuse and $h({\mathcal{A}}_{i}:{\mathcal{M}})=0$, then ${\mathcal{A}}_{i}$ must be contained in ${\mathcal{M}}_{i}$. Indeed, from Theorem F, we see that ${\mathcal{M}}_{1}\vee{\mathcal{A}}_{1}$ is freely independent from $\mathcal{M}_{2}$ and this in particular means that ${\mathcal{M}}_{1}\vee{\mathcal{A}}_{1}={\mathcal{M}}_{1}$. In Theorem F (1), we do not assume ${\mathcal{A}}_{i}$ is in ${\mathcal{M}}$, only in ${\mathcal{M}}^{{\mathcal{U}}}$, but we still obtain free independence of ${\mathcal{M}}_{i}\vee{\mathcal{A}}_{i}$. However, in this more general setting, one cannot conclude that ${\mathcal{A}}_{i}\subseteq{\mathcal{M}}_{i}^{{\mathcal{U}}}\subseteq{\mathcal{M}}^{{\mathcal{U}}}$. For instance, suppose that ${\mathcal{M}}_{i}^{{\mathcal{U}}}$ is commutative; then let ${\mathcal{A}}_{i}$ be a copy of the hyperfinite $\mathrm{II}_{1}$ factor that intersects ${\mathcal{M}}_{i}$ diffusely, which must exist since all Haar unitaries in ${\mathcal{M}}^{{\mathcal{U}}}$ are conjugate. Then ${\mathcal{A}}_{i}$ cannot be contained in the commutative algebra ${\mathcal{M}}_{i}^{{\mathcal{U}}}$.
1.2. Broader context and motivation
The broader motivation for our work includes a long history of results about the von Neumann algebra $L(F_{\infty})$, going back to Murray and von Neumann’s first papers in the subject [43, §6]. The challenge of understanding its structure has been addressed with an astonishing diversity of tools, including not only the probabilistic methods that our paper draws on [62, 21, 37, 27, 28], but also deformation rigidity theory [53, 55, 45, 57]; amenable actions, $C^{*}$ and boundary theory [44, 12, 13]; closable derivations [47, 48]; free harmonic analysis and non-commutative $L_{p}$ space theory [42], which was a key ingredient in Ioana and Houdayer’s result [32]. We also point out that Popa showed the existence of elements in an ultraproduct freely independent from many subalgebras using incremental patching techniques [54, 56].
We were particularly motivated by the question of elementary equivalence of von Neumann algebras, and in particular of $L(F_{\infty})$ and matrix ultraproducts. The introduction of ultraproducts in functional analysis naturally inspired the classification question of when ${\mathcal{M}}^{{\mathcal{U}}}$ and ${\mathcal{N}}^{{\mathcal{V}}}$ are isomorphic for various ultrafilters ${\mathcal{U}}$ and ${\mathcal{V}}$. Using mathematical logic, isomorphism of ${\mathcal{M}}^{{\mathcal{U}}}$ and ${\mathcal{M}}^{{\mathcal{V}}}$ for different ultrapowers depends on the continuum hypothesis [15, Proposition 3.3]. But model theory also provides a powerful tool, known as the Keisler–Shelah theorem, which says that ${\mathcal{M}}$ and ${\mathcal{N}}$ admit some isomorphic ultrapowers if and only if they have the same first-order theory (see [17, §2.2]), which here we understand in the sense of model theory for metric structures [5]. In this case, ${\mathcal{M}}$ and ${\mathcal{N}}$ are said to be elementarily equivalent.
The classification of tracial von Neumann algebras up to elementary equivalence is a challenging problem [23, §4] and in particular very little is known in the setting of $\mathrm{II}_{1}$ factors without property Gamma. For instance, we do not know if the matrix ultraproducts with different ultrafilters are elementarily equivalent [17, §5] [35, §5.2], we do not know if $L(F_{m})$ and $L(F_{n})$ are elementarily equivalent for $m\neq n$ [24], nor do we know if $L(F_{\infty})$ is elementarily equivalent to a matrix ultraproduct [23, Question 4.6]. One could try to distinguish the theories of these algebras with certain first-order sentences relating to familiar properties such as commutation, free independence, and the like. For instance, the construction in [10] and with minor modifications in [32, Theorems F and G] and [38, §2.4] of non-Gamma $\mathrm{II}_{1}$ factors that are not elementarily equivalent is motivated by statements such as ‘‘for all unitaries $u_{1}$ and $u_{2}$ with $u_{1}^{2}=u_{2}^{3}=1$, there (approximately) exist Haar unitaries $v_{1}$ and $v_{2}$ with $[u_{1},v_{1}]=[v_{1},v_{2}]=[v_{2},u_{2}]=0$,’’ see [10, Remark 5.9]; these results fit into the formalism of sequential commutation introduced in [38]. Houdayer and Ioana’s theorem on freeness of commutants in [32] implies that ‘‘there exist unitaries $u_{1}$, $u_{2}$, …such that for all $b_{1}$, $b_{2}$, …with $[u_{j},b_{j}]=0$ (approximately), we have that $b_{1}$, $b_{2}$are (approximately) freely independent’’ holds in $L(F_{\infty})$, while our first result Theorem B shows that a similar statement holds in matrix ultraproducts; see Remark 4.5 for a more precise statement.
Moreover, Theorem C yields further first-order statements that hold in matrix ultraproducts; see Lemma 4.2. Since $u_{1}$, $u_{2}$, …arise from Haar random unitaries in the large-$n$ limit, our results also give some information about the values of first-order formulas on independent random unitaries. While Voiculescu’s asymptotic freeness theory [60, 63] describes the *-moments of random matrices, precious little information is known about formulas that involve quantifiers; for related ideas, see [35, §5.2], [18, §4].
1.3. Notation and Organization
Here tracial von Neumann algebra refers to a von Neumann algebra ${\mathcal{M}}$ with a fixed faithful normal tracial state $\operatorname{tr}^{{\mathcal{M}}}$. We write $\lVert x\rVert_{2}=\operatorname{tr}^{{\mathcal{M}}}(x^{*}x)^{1/2}$. In particular, when ${\mathcal{M}}=M_{n}({\mathbb{C}})$, we denote the normalized trace by $\operatorname{tr}_{n}$ and write $\lVert x\rVert_{2}=\operatorname{tr}_{n}(x^{*}x)^{1/2}$. We assume that familiarity with basic theory of tracial von Neumann algebras (see for instance [58, 65, 6, 1]) as well as ultraproducts of tracial von Neumann algebras (see for instance [8, Appendix A], [1, §5.4]).
Concerning the organization of the paper, we first give a self-contained proof of Theorems A and B through concentration and volumetric analysis in §2. Then we prepare ingredients about model theory and $1$-bounded entropy in §3, which are subsequentily used in the proof of Theorems C and F in §4.
Acknowledgements
We are indebted to Ben Hayes for encouragement, generosity, and insightful discussions. We thank Adrian Ioana for bringing to our attention the problem of free independence of commutants in matrix ultraproducts. DJ thanks Jonathan Shi and Juspreet Singh for sharing a question about an application of random matrices in computer science that motivated uniform asymptotic freeness. We thank Greg Patchell and David Gao for various helpful discussions. We thank Yoonje Jeong as well as the anonymous referees for comments that improved the correctness and clarity of the paper.
2. Freeness of commutants
In this section, after recalling some elementary facts about Haar random unitaries and diagonalization, we give our concentration of measure argument for uniform asymptotic freeness, and finally prove Theorems A and B.
2.1. Useful facts about diagonalization and approximate commutants
In order to analyze the commutant of $u_{j}$ more easily, it will be convenient to diagonalize the Haar random unitary $U_{j}^{(n)}$. In fact, up to a small error, we will be able to arrange that $U_{j}^{(n)}$ has the form $V_{j}^{(n)}A^{(n)}(V_{j}^{(n)})^{*}$ for deterministic $A^{(n)}$ with evenly spaced eigenvalues. Thus, to prove the main results, we can perform our analysis on the model given by conjugates of a diagonal matrix, and then transfer them over to $U_{j}^{(n)}$’s.
The following are useful elementary observations that can be considered folklore in random matrix theory, but we include the proofs here for completeness. Some of the lemmas in this section will actually be used several times.
Lemma 2.1.
For $k\in{\mathbb{N}}$, let $[k]=\{1,\dots,k\}$. For $j\in[k]$, let $I_{k,j}$ be the interval $[2\pi(j-1)/k,2\pi j/k)$ viewed as a subset of the unit circle $S^{1}$. Let
$\mathcal{O}_{k}=\{\mu\in\mathcal{P}(S^{1}):~{}\forall j\in[k],~{}1/k-1/k^{2}<\mu(I_{k,j}^{\circ})\leq\mu(\overline{I}_{k,j})<1/k+1/k^{2}\}.$
Then $\mathcal{O}_{k}$ is an open neighborhood in $\mathcal{P}(S^{1})$ of the Haar measure.
Let $A^{(n)}=\operatorname{diag}(1,\zeta_{n},\zeta_{n}^{2},\dots,\zeta_{n}^{n-1})$, where $\zeta_{n}=e^{2\pi i/n}$. Let $B$ be a diagonal matrix with eigenvalues $e^{i\lambda_{1}}$, …, $e^{i\lambda_{n}}$ where $0\leq\lambda_{1}\leq\dots\leq\lambda_{n}<2\pi$. If the empirical spectral distribution of $B$ is in $\mathcal{O}_{k}$, then $\lVert A^{(n)}-B\rVert<4\pi/k$.
Proof.
[…]
For the second claim, let $\mu$ be the empirical spectral distribution of $B$, and suppose $\mu\in\mathcal{O}_{k}$. Then for $j\in[k]$,
$(j-1)/k\leq j/k-j/k^{2}\leq\mu(I_{k,1}\cup\dots\cup I_{k,j})\leq j/k+j/k^{2}\leq(j+1)/k.$
Therefore, if $t\in[n]$ with $t/n>(j+1)/k$, then $\lambda_{t}^{(n)}\geq 2\pi j/k$, and if $t\in[n]$ with $t/n<(j-1)/k$, then $\lambda_{t}^{(n)}\leq 2\pi j/k$. Overall, this implies that
$|\lambda_{t}^{(n)}-2\pi t/n|\leq\frac{4\pi}{k},\text{ hence }|\zeta_{t}^{(n)}-e^{i\lambda_{t}^{(n)}}|\leq\frac{4\pi}{k}$
since the complex exponential function is $1$-Lipschitz. Therefore, $\lVert B-A^{(n)}\rVert\leq 4\pi/k$ as desired. ∎
Proposition 2.2.
Let $A^{(n)}=\operatorname{diag}(1,\zeta_{n},\zeta_{n}^{2},\dots,\zeta_{n}^{n-1})$, where $\zeta_{n}=e^{2\pi i/n}$. Then there exists a family of independent Haar random unitaries $U_{1}^{(n)}$, …, $U_{m}^{(n)}$ and another family of independent Haar random unitaries $V_{1}^{(n)}$, …, $V_{m}^{(n)}$ on the same probability space $(\Omega,\mathcal{F},P)$ such that
$\lim_{n\to\infty}\sum_{j=1}^{m}\lVert U_{j}^{(n)}-V_{j}^{(n)}A^{(n)}(V_{j}^{(n)})^{*}\rVert=0\text{ almost surely.}$.
Remark 2.3.
The conclusion in the proposition shows that the operator norm goes to zero. For this paper, we only need the weaker statement that the $2$-norm goes to zero.
Proof of Proposition 2.2.
First, we remark the following: If $X$ is an $n\times n$ Haar random unitary and $Y$ is an $n\times n$ random unitary independent of $X$, then $XY$ and $YXY^{*}$ are Haar random unitaries. In the case where $Y$ is deterministic, $YX$ is a Haar unitary since the Haar measure is left-invariant, and then $YXY^{*}$ is a Haar unitary also since the Haar measure is right-invariant. Now consider $Y$ that is random and independent of $X$. Since $(X,Y)$ are independent, the joint distribution $(X,Y)$ has a disintegration given by conditioning on the value of $Y$. Moreover, the conditional distributions of $XY$ and $YXY^{*}$ given that $Y$ is some fixed value $y$ are the Haar measure on the unitary group, which follows from the case of deterministic $Y$ handled above. Now the distribution of $YX$ and $YXY^{*}$ respectively are obtained by integrating the conditional distributions given $Y=y$ with respect to the marginal distribution of $Y$. Hence, they are also equal to the Haar measure on the unitary group.
Now let $X_{1}^{(n)}$, …, $X_{m}^{(n)}$, $Y_{1}^{(n)}$, …, $Y_{m}^{(n)}$ be independent Haar random unitary matrices. Let $U_{j}^{(n)}=Y_{j}^{(n)}X_{j}^{(n)}(Y_{j}^{(n)})^{*}$. By the foregoing argument, $U_{1}^{(n)}$, …, $U_{m}^{(n)}$ are independent Haar random unitaries. By the spectral theorem, we may write $X_{j}^{(n)}=W_{j}^{(n)}B^{(n)}(W_{j}^{(n)})^{*}$, where $B^{(n)}$ is diagonal and $W_{j}^{(n)}$ is unitary. For $\zeta$ on the unit circle, let $\arg(\zeta)$ be the value of the argument that is in $[0,2\pi)$. If $X_{j}^{(n)}$ has distinct eigenvalues, then there is a unique choice of $B_{j}^{(n)}$ where the arguments of the diagonal entries are in increasing order. Moreover, in the case when $X_{j}^{(n)}$ has distinct eigenvalues (which happens almost surely), the choice of $W_{j}^{(n)}$ such that $X_{j}^{(n)}=W_{j}^{(n)}B^{(n)}(W_{j}^{(n)})^{*}$ is unique as well. It is also straightforward to check that it depends on $X_{j}^{(n)}$ in a Borel-measurable manner. Since the $X_{j}^{(n)}$’s are independent of the $Y_{j}^{(n)}$’s, we also have that $W_{j}^{(n)}$’s are independent of the $Y_{j}^{(n)}$’s, and therefore $V_{j}^{(n)}=Y_{j}^{(n)}W_{j}^{(n)}$ is a Haar random unitary, and of course $V_{1}^{(n)}$, …, $V_{m}^{(n)}$ are independent since $V_{j}^{(n)}$ only depends on $X_{j}^{(n)}$ and $Y_{j}^{(n)}$ for each $j$. Overall, we have
$U_{j}^{(n)}=V_{j}^{(n)}B_{j}^{(n)}(V_{j}^{(n)})^{*},$
where $U_{1}^{(n)}$, …, $U_{m}^{(n)}$ are independent Haar random unitaries, $V_{1}^{(n)}$, …, $V_{m}^{(n)}$ are independent Haar random unitaries (though not independent of $U_{j}^{(n)}$), and $B_{j}^{(n)}$ is diagonal with eigenvalues listed in order of increasing argument in $[0,2\pi)$.
Note that
$\lVert U_{j}^{(n)}-V_{j}^{(n)}A_{j}^{(n)}(V_{j}^{(n)})^{*}\rVert_{2}=\lVert V_{j}^{(n)}B_{j}^{(n)}(V_{j}^{(n)})^{*}-V_{j}^{(n)}A_{j}^{(n)}(V_{j}^{(n)})^{*}\rVert_{2}=\lVert B_{j}^{(n)}-A_{j}^{(n)}\rVert_{2}.$
Therefore, to complete the proof, it suffices to show that $\lim_{n\to\infty}\lVert B_{j}^{(n)}-A_{j}^{(n)}\rVert_{2}=0$ almost surely.
Let $\mu_{j}^{(n)}$ be the empirical spectral distribution of $U^{(n)}$, that is, the (random) probability measure on the circle that has a point mass of $1/n$ at each eigenvalue of $U_{j}^{(n)}$. Note that empirical spectral distributions of $U_{j}^{(n)}$ and $B_{j}^{(n)}$ are the same. By standard results about random unitary matrices (see for instance [40, Theorem 4.13]), we have that almost surely $\mu_{j}^{(n)}$ converges weak-* to the Haar measure on the unit circle. Hence, almost surely, $\mu_{j}^{(n)}$ is eventually in the neighborhood $\mathcal{O}_{k}$ in Lemma 2.1. It follows that $\lVert B_{j}^{(n)}-A^{(n)}\rVert$ is eventually less than $4\pi/k$, and since $k$ is arbitrary, this completes the proof. ∎
This result on diagonalization enables us to reduce the study of approximate commutants of $U_{j}^{(n)}$ to the case of $A^{(n)}$. In this setting, the approximate commutant is described by $\epsilon$-diagonal or band matrices.
Notation 2.4.
For $i,j\in[n]$, let $d_{n}(i,j)$ denote the distance of $i$ and $j$ modulo $n$. Let
$\mathcal{D}_{\epsilon}^{(n)}:=\{B\in M_{n}({\mathbb{C}}):B_{i,j}=0\text{ when }d_{n}(i,j)>\epsilon n\}.$
Lemma 2.5.
Let $A^{(n)}=\operatorname{diag}(1,\zeta_{n},\zeta_{n}^{2},\dots,\zeta_{n}^{n-1})$ where $\zeta_{n}=e^{2\pi i/n}$. Let $B\in M_{n}(\mathbb{C})$. Then for every $\epsilon>0$, there exists $B_{\epsilon}\in\mathcal{D}_{\epsilon}^{(n)}$ with
$\lVert B-B_{\epsilon}\rVert_{2}\leq\frac{1}{\epsilon}\lVert[A^{(n)},B]\rVert_{2},\qquad\lVert B_{\epsilon}\rVert\leq 3\lVert B\rVert.$
Proof.
Fix $\epsilon$ and $B$. In the case where $\epsilon<2/n$, we take $B_{\epsilon}$ to be the projection of $B$ onto diagonal matrices. Note that
$\lVert B-B_{\epsilon}\rVert_{2}^{2}=\frac{1}{n}\sum_{j\neq k}|B_{j,k}|^{2},$
while
(2.1) $\lVert[A^{(n)},B]\rVert_{2}^{2}=\frac{1}{n}\sum_{j\neq k}|\zeta_{n}^{j}-\zeta_{n}^{k}|^{2}|B_{j,k}|^{2}.$
Note
$|\zeta_{n}^{j}-\zeta_{n}^{k}|^{2}=|1-\zeta_{n}^{j-k}|^{2}=2-2\cos(2\pi d_{n}(j,k)/n)\geq 2-2\cos(2\pi/n)$
Note that for $x\in[-\pi,\pi]$, we have $1-\cos x=2\sin(x/2)^{2}\geq 2x^{2}/\pi^{2}$ since $|\sin(x/2)|\geq|x|/\pi$ on $[-\pi,\pi]$ using concavity of the sine function. Hence,
$|\zeta_{n}^{j}-\zeta_{n}^{k}|^{2}\geq\frac{4}{\pi^{2}}\left(\frac{2\pi}{n}\right)^{2}=\frac{16}{n^{2}}\geq 4\epsilon^{2}.$
Overall,
$\lVert B-B_{\epsilon}\rVert_{2}\leq\frac{1}{2\epsilon}\lVert[A^{(n)},B]\rVert_{2}$
Now suppose that $\epsilon\geq 2/n$. Let $m=\lfloor n\epsilon/2\rfloor\geq 2$. Write $n=qm+r$ for some $r\in\{0,\dots,m\}$. For $j=1$, …, $q$, let $P_{j}$ be the projection onto the basis vectors $e_{(j-1)m+1}$, …, $e_{jm}$; and let $P_{q+1}$ be the projection onto the last $r$ basis vectors. Let
$B_{\epsilon}=\sum_{d_{q+1}(j,k)\leq 1}P_{j}BP_{k}.$
Note that all the indices $(j^{\prime},k^{\prime})$ where $(B_{\epsilon})_{j^{\prime},k^{\prime}}\neq 0$ occur when $d_{n}(j^{\prime},k^{\prime})\leq 2m\leq n\epsilon$, and thus $B_{\epsilon}\in\mathcal{D}_{\epsilon}^{(n)}$. Moreover, by writing
$B_{\epsilon}=\sum_{j=1}^{q+1}P_{j}BP_{j}+\sum_{j=1}^{q+1}P_{j}BP_{j+1}+\sum_{j=1}^{q+1}P_{j}BP_{j-1}$
(indices considered modulo $q+1$), we see that $\lVert B_{\epsilon}\rVert\leq 3\lVert B\rVert$. Next, note that the indices $(j^{\prime},k^{\prime})$ where $(B-B_{\epsilon})_{j^{\prime},k^{\prime}}\neq 0$ must satisfy
$d_{n}(j^{\prime},k^{\prime})\geq m\geq n\epsilon/2-1\geq n\epsilon/4,$
and so
$2-2\cos(2\pi d_{n}(j^{\prime},k^{\prime})/n)\geq 2-2\cos(\pi\epsilon/2)\geq\frac{4}{\pi^{2}}\left(\frac{\pi\epsilon}{2}\right)^{2}=\epsilon^{2}.$
Therefore,
$\lVert B-B_{\epsilon}\rVert_{2}^{2}\leq\frac{1}{n}\sum_{d_{n}(j,k)>n\epsilon/4}|B_{j,k}|^{2}\leq\frac{1}{\epsilon^{2}}\frac{1}{n}\sum_{j\neq k}|\zeta_{n}^{j}-\zeta_{n}^{k}|^{2}|B_{j,k}|^{2}\leq\frac{1}{\epsilon^{2}}\lVert[A^{(n)},B]\rVert_{2}^{2}.$
Corollary 2.6.
Fix a free ultrafilter ${\mathcal{U}}$ on ${\mathbb{N}}$ and write ${\mathcal{Q}}=\prod_{n\to{\mathcal{U}}}M_{n}({\mathbb{C}})$. Let $A^{(n)}=\operatorname{diag}(1,\zeta_{n},\zeta_{n}^{2},\dots,\zeta_{n}^{n-1})$, and let $a=[A^{(n)}]_{n\in{\mathbb{N}}}\in{\mathcal{Q}}$. Let $V_{j}^{(n)}$ for $j\in{\mathbb{N}}$ be independent Haar random unitaries. Fix an outcome $\omega$ and let $v_{j}(\omega)=[V_{j}^{(n)}(\omega)]\in{\mathcal{Q}}$. Let
${\mathcal{D}}_{\epsilon}=\{[X^{(n)}]_{n\in{\mathbb{N}}}\in{\mathcal{Q}}:X^{(n)}\in\mathcal{D}_{\epsilon}^{(n)}\text{ for }n\in{\mathbb{N}}\}.$
Then
$\{v_{j}(\omega)av_{j}(\omega)^{*}\}^{\prime}\cap{\mathcal{Q}}=\bigcap_{\epsilon>0}v_{j}(\omega)\mathcal{D}_{\epsilon}v_{j}(\omega)^{*}$
Proof.
Since commutants respect conjugation, it suffices to show that $\{a\}^{\prime}\cap{\mathcal{Q}}=\bigcap_{\epsilon>0}\mathcal{D}_{\epsilon}$. Let $b=[B^{(n)}]_{n\in{\mathbb{N}}}$ be an element commuting with $a$, and fix $\epsilon>0$. Let $B_{\epsilon}^{(n)}$ be as in Lemma 2.5. Since $\lVert[A^{(n)},B^{(n)}]\rVert_{2}\to 0$ along the ultrafilter, we obtain that $\lVert B^{(n)}-B_{\epsilon}^{(n)}\rVert_{2}\to 0$ as well. Hence, $b=[B_{\epsilon}^{(n)}]_{n\in{\mathbb{N}}}$ is in $\mathcal{D}_{\epsilon}$, as desired. ∎
2.2. Uniform asymptotic freeness via concentration of measure
For Theorems B and A, we will proceed as follows. By Proposition 2.2, we can replace $U_{j}^{(n)}$ with $V_{j}^{(n)}A^{(n)}(V_{j}^{(n)})^{*}$. In this section, we will argue that $V_{j}^{(n)}\mathcal{D}_{\epsilon}^{(n)}(V_{j}^{(n)})^{*}$ are asymptotically free as $n\to\infty$ up to some error tolerance depending on $\epsilon$. Then since matrices that approximately commute with $V_{j}^{(n)}A^{(n)}(V_{j}^{(n)})^{*}$ will, for any $\epsilon$, be approximately in $V_{j}^{(n)}\mathcal{D}_{\epsilon}^{(n)}(V_{j}^{(n)})^{*}$, we will obtain the desired results.
Hence, our present goal is to obtain a uniform approximate asymptotic freeness result for $V_{j}^{(n)}\mathcal{D}_{\epsilon}^{(n)}(V_{j}^{(n)})^{*}$. This is based on Voiculescu’s famous asymptotic freeness theorem.
Theorem 2.7 (Voiculescu’s asymptotic freeness [63]).
Let $U_{1}^{(n)}$, …, $U_{m}^{(n)}$ be $n\times n$ Haar random unitaries. Let $i_{1}\neq i_{2}\neq\dots i_{k}$. Then
$\lim_{n\to\infty}\sup_{X_{1},\dots,X_{k}\in B_{1}^{M_{n}({\mathbb{C}})}}|\mathbb{E}\operatorname{tr}_{n}\left[U_{i_{1}}^{(n)}(X_{1}-\operatorname{tr}_{n}(X_{1}))(U_{i_{1}}^{(n)})^{*}\dots U_{i_{k}}^{(n)}(X_{k}-\operatorname{tr}_{n}(X_{k}))(U_{i_{k}}^{(n)})^{*}\right]|=0.$
Proof.
Let $B_{r}^{M_{n}({\mathbb{C}})}$ denote the $r$-ball with respect to operator norm. For each $n$, fix $X_{1}^{(n)}$, …, $X_{k}^{(n)}$ in $B_{1}^{M_{n}({\mathbb{C}})}$ which maximize
(2.2) $|\mathbb{E}\operatorname{tr}_{n}\left[U_{i_{1}}^{(n)}(X_{1}^{(n)}-\operatorname{tr}_{n}(X_{1}^{(n)}))(U_{i_{1}}^{(n)})^{*}\dots U_{i_{k}}^{(n)}(X_{k}^{(n)}-\operatorname{tr}_{n}(X_{k}^{(n)}))(U_{i_{k}}^{(n)})^{*}\right]|.$
A maximizer exists because the quantity depends continuously on the $X_{j}$’s. It follows from [63, Corollary 2.5] that (2.2) converges to 00 as $n\to\infty$. ∎
In order to apply this result uniformly to all the matrices $X_{j}\in\mathcal{D}_{\epsilon}^{(n)}$, we rely on the high-dimensional concentration of measure phenomenon for Haar unitaries, which has been a staple of random matrix theory since [4, 25]. We recall that the unitary group $\mathbb{U}_{n}$ (equipped with the Riemannian metric associated to the inner product $\langle\cdot,\cdot\rangle_{\operatorname{Tr}_{n}}$) satisfies the log-Sobolev inequality with constant $6/n$ [41, Theorem 15]. One can easily deduce the log-Sobolev inequality for the product of several copies of $\mathbb{U}_{n}$; see e.g. [39, Corollary 5.7], [40, Theorem 5.9]. This in turn implies that it satisfies the Herbst concentration estimate; see e.g. [2, Lemma 2.3.3], [40, Theorem 5.5]. After renormalizing the metric to $\langle\cdot,\cdot\rangle_{\operatorname{tr}_{n}}$, one obtains the following concentration bound. See [31, §5.3] for further explanation.
Lemma 2.8 (Concentration).
Let $U_{1}^{(n)}$, …, $U_{m}^{(n)}$ be $n\times n$ Haar random unitaries, and let $f:\mathbb{U}_{n}^{m}\to\mathbb{C}$ be Lipschitz with respect to $\lVert\cdot\rVert_{2}$. Then
$P(|f(U_{1}^{(n)},\dots,U_{m}^{(n)})-\mathbb{E}[f(U_{1}^{(n)},\dots,U_{m}^{(n)})]|\geq\delta)\leq 4e^{-n^{2}\delta^{2}/12\lVert f\rVert_{\operatorname{Lip}}^{2}}$
Concentration allows us to deduce the following uniform asymptotic freeness result, which is our main technical tool. Here we replace $\mathcal{D}_{\epsilon}^{(n)}$ with a more general set $S_{j}^{(n)}$ that has relatively small covering numbers, which also depends on the index $j$ representing the position in the string or monomial; this added generality will be used in the proof of Lemma 4.1.
Notation 2.9.
Let $S$ be a subset of a metric space $X$. Then $K_{\epsilon}(S)$ is defined as the smallest cardinality of a set $\Omega\subseteq X$ such that the $\epsilon$-neighborhood of $\Omega$ covers $S$.
Lemma 2.10 (Uniform asymptotic freeness).
Fix $k\in{\mathbb{N}}$. For each $n\in{\mathbb{N}}$ and $j\in[k]$, let $S_{j}^{(n)}\subseteq B_{1}^{M_{n}({\mathbb{C}})}$ and suppose that
(2.3) $\lim_{n\to{\mathcal{U}}}\frac{1}{n^{2}}\log K_{\epsilon}(S_{j}^{(n)})<\delta.$
Let $i_{1}\neq i_{2}\neq\dots i_{k}$ in ${\mathbb{N}}$. Then almost surely
(2.4) $\lim_{n\to{\mathcal{U}}}\sup_{X_{1}\in S_{1}^{(n)}}\dots\sup_{X_{k}\in S_{k}^{(n)}}\left|\operatorname{tr}_{n}\left[U_{i_{1}}^{(n)}(X_{1}-\operatorname{tr}_{n}(X_{1}))(U_{i_{1}}^{(n)})^{*}\dots U_{i_{k}}^{(n)}(X_{k}-\operatorname{tr}_{n}(X_{k}))(U_{i_{k}}^{(n)})^{*}\right]\right|\\ \leq 4k\epsilon+2k\sqrt{12k\delta}.$
The same statement also holds when $\lim_{n\to{\mathcal{U}}}$ is replaced by $\limsup_{n\to\infty}$ in both the hypothesis and the conclusion.
Proof.
First, fix a set $\Omega_{j}^{(n)}$ such that the $\epsilon$-neighborhood of $\Omega_{j}^{(n)}$ covers $S_{j}^{(n)}$, and such that $|\Omega_{j}^{(n)}|=K_{\epsilon}^{(n)}(S_{j}^{(n)})$. Although $\Omega_{j}^{(n)}$ is initially not assumed to be a subset of $S_{j}^{(n)}$, we can replace each element of $\Omega_{j}^{(n)}$ with an element from $S_{j}^{(n)}$ in its $\epsilon$-ball. Hence, WLOG assume that $\Omega_{j}^{(n)}\subseteq S_{j}^{(n)}$ such that $S_{j}^{(n)}$ is in the $2\epsilon$-neighborhood of $\Omega_{j}^{(n)}$.
Let $m=\max(i_{1},\dots,i_{k})$. Let
$f(X_{1},\dots,X_{k},U_{1}^{(n)},\dots,U_{m}^{(n)})=\operatorname{tr}_{n}\left[U_{i_{1}}^{(n)}(X_{1}-\operatorname{tr}_{n}(X_{1}))(U_{i_{1}}^{(n)})^{*}\dots U_{i_{k}}^{(n)}(X_{k}-\operatorname{tr}_{n}(X_{k}))(U_{i_{k}}^{(n)})^{*}\right]$
Since $\lVert X_{j}-\operatorname{tr}_{n}(X_{j})\rVert\leq\lVert X_{j}\rVert\leq 1$, we see that $f$ is a $2k$-Lipschitz function of $(U_{1}^{(n)},\dots,U_{m}^{(n)})$ with respect to $\lVert\cdot\rVert_{2}$. In particular,
(2.5) $\sup_{X_{1}\in S_{1}^{(n)}}\dots\sup_{X_{k}\in S_{k}^{(n)}}|f(X_{1},\dots,X_{k},U_{1}^{(n)},\dots,U_{m}^{(n)})|\\ \leq 4k\epsilon+\sup_{X_{1}\in\Omega_{1}^{(n)}}\dots\sup_{X_{k}\in\Omega_{k}^{(n)}}|f(X_{1},\dots,X_{k},U_{1}^{(n)},\dots,U_{m}^{(n)})|.$
Now by Lemma 2.8, we have for each $(X_{1},\dots,X_{k})$ that
(2.6) $P(|f(X_{1},\dots,X_{k},U_{1}^{(n)},\dots,U_{m}^{(n)})-\mathbb{E}[f(X_{1},\dots,X_{k},U_{1}^{(n)},\dots,U_{m}^{(n)})]|\geq 2k\sqrt{12k\delta})\\ \leq 4e^{-n^{2}12k\delta(2k)^{2}/12(2k)^{2}}=4e^{-n^{2}k\delta}$
Hence, by a union bound,
$P\left(\sup_{X_{1}\in\Omega_{1}^{(n)}}\dots\sup_{X_{k}\in\Omega_{k}^{(n)}}|f(X_{1},\dots,X_{k},U_{1}^{(n)},\dots,U_{m}^{(n)})-\mathbb{E}[f(X_{1},\dots,X_{k},U_{1}^{(n)},\dots,U_{m}^{(n)})]|\geq 2k\sqrt{12k\delta}\right)\\ \leq 4e^{-n^{2}k\delta}\prod_{j=1}^{k}K_{\epsilon}(S_{j}^{(n)}).$
By (2.3), there exists $\delta^{\prime}<\delta$ and some $A\in{\mathcal{U}}$ such that for all $n\in A$, for $j=1$, …, $k$, we have $K_{\epsilon}(S_{j}^{(n)})\leq e^{n^{2}\delta^{\prime}}$. In particular, the previous equation is bounded by
$4e^{-n^{2}k\delta}e^{n^{2}k\delta^{\prime}}=e^{-n^{2}k(\delta-\delta^{\prime})}.$
Since $\delta^{\prime}<\delta$, this quantity is summable over $n\in{\mathbb{N}}$, and in particular, it is summable over $n\in A$. Hence, by the Borel-Cantelli lemma, almost surely
(2.7) $\lim_{\begin{subarray}{c}n\to\infty\\ n\in A\end{subarray}}\sup_{X_{1}\in\Omega_{1}^{(n)}}\dots\sup_{X_{k}\in\Omega_{k}^{(n)}}|f(X_{1},\dots,X_{k},U_{1}^{(n)},\dots,U_{m}^{(n)})-\mathbb{E}[f(X_{1},\dots,X_{k},U_{1}^{(n)},\dots,U_{m}^{(n)})]|\\ \leq 2k\sqrt{12k\delta}.$
In particular, this holds for the limit as $n\to{\mathcal{U}}$. Moreover, by Theorem 2.7, we have that
(2.8) $\lim_{n\to\infty}\sup_{X_{1}\in\Omega_{1}^{(n)}}\dots\sup_{X_{k}\in\Omega_{k}^{(n)}}|\mathbb{E}[f(X_{1},\dots,X_{k},U_{1}^{(n)},\dots,U_{m}^{(n)})]|=0.$
[…]
Finally, we record the following standard estimate on the covering number of $\mathcal{D}_{\epsilon}^{(n)}$.
Lemma 2.11.
For $0<\epsilon<R$, we have
$\frac{1}{n^{2}}\log K_{\epsilon}(\{X\in\mathcal{D}_{\epsilon}^{(n)}:\lVert X\rVert_{2}\leq R\})\leq 2\epsilon\log\frac{3R}{\epsilon}.$
Proof.
Note that $\dim_{{\mathbb{R}}}\mathcal{D}_{\epsilon}^{(n)}\leq 2\epsilon n^{2}$. Let $\Omega$ be a maximal set of $\epsilon$-separated points in $S=\{X\in\mathcal{D}_{\epsilon}^{(n)}:\lVert X\rVert_{2}\leq R\}$. Then the $\epsilon$-neighborhood of $\Omega$ covers $S$ by maximality. The $\epsilon/2$ balls with centers in $\Omega$ are disjoint, and they are contained in the $R+\epsilon/2\leq 3R/2$-ball centered at zero. Therefore, $|\Omega|$ is at most
$\left(\frac{3R/2}{\epsilon/2}\right)^{2\epsilon n^{2}}=\left(\frac{3R}{\epsilon}\right)^{2\epsilon n^{2}}.\qed$
2.3. Proof of freeness of commutants
All the pieces are now in place to prove Theorems A and B. We start with Theorem B first since it involves fewer approximation arguments.
Proof of Theorem B.
As in Proposition 2.2, let $U_{j}^{(n)}$ be independent Haar random unitaries. Let $A^{(n)}=\operatorname{diag}(1,\zeta_{n},\dots,\zeta_{n}^{n-1}$, where $\zeta_{n}=e^{2\pi i/n}$, and let $V_{j}^{(n)}$ be independent Haar random unitaries, such that $\lVert U_{j}^{(n)}-V_{j}^{(n)}A^{(n)}(V_{j}^{(n)})^{*}\rVert_{2}\to 0$ almost surely as $n\to\infty$. Let $u_{j}(\omega)=[U_{j}^{(n)}(\omega)]_{n\in{\mathbb{N}}}\in{\mathcal{Q}}$, and let $v_{j}(\omega)=[V_{j}^{(n)}(\omega)]_{n\in{\mathbb{N}}}\in{\mathcal{Q}}$, and let $a=[A^{(n)}]_{n\in{\mathbb{N}}}\in{\mathcal{Q}}$. Thus, almost surely $u_{j}(\omega)=v_{j}(\omega)av_{j}(\omega)^{*}$, and so
$\{u_{j}(\omega)\}^{\prime}\cap{\mathcal{Q}}=\{v_{j}(\omega)av_{j}(\omega)^{*}\}^{\prime}\cap{\mathcal{Q}}=\bigcap_{\epsilon>0}v_{j}(\omega)\mathcal{D}_{\epsilon}v_{j}(\omega)^{*},$
where the second equality follows from Corollary 2.1 with $\mathcal{D}_{\epsilon}=\{[X^{(n)}]_{n\in{\mathbb{N}}}\in{\mathcal{Q}}:X^{(n)}\in{\mathcal{D}}_{\epsilon}^{(n)}\}$.
For each word $i_{1}\neq\dots\neq i_{k}$, and for each $m\in{\mathbb{N}}$, taking $\epsilon=1/m$ in Lemma 2.11 and $\delta=2\epsilon\log(3R/\epsilon)$ in Lemma 2.10, we have that almost surely
(2.9) $\limsup_{n\to\infty}\sup_{X_{1}\in\mathcal{D}_{1/m}^{(n)}}\dots\sup_{X_{k}\in\mathcal{D}_{1/m}^{(n)}}\left|\operatorname{tr}_{n}\left[V_{i_{1}}^{(n)}(X_{1}-\operatorname{tr}_{n}(X_{1}))(V_{i_{1}}^{(n)})^{*}\dots V_{i_{k}}^{(n)}(X_{k}-\operatorname{tr}_{n}(X_{k}))(V_{i_{k}}^{(n)})^{*}\right]\right|$
(2.10) $\leq 4k/m+2k\sqrt{24k\log(3Rm)/m}.$
There are only countably many words and values of $m$, so almost surely this holds for all $i_{1}$, …, $i_{k}$ and $m$, and also $\lVert U_{j}^{(n)}-V_{j}^{(n)}A^{(n)}(V_{j}^{(n)})^{*}\rVert_{2}\to 0$. In the rest of the proof, we fix an outcome $\omega$ in this almost sure event.
To show free independence of the commutants, fix $k\in{\mathbb{N}}$, fix a word $i_{1}\neq i_{2}\neq\dots\neq i_{k}$, and for $j\in[k]$, fix $y_{j}\in\{u_{i_{j}}(\omega)\}^{\prime}\cap{\mathcal{Q}}$ with $\operatorname{tr}^{{\mathcal{Q}}}(y_{j})=0$. Let $x_{j}=v_{i_{j}}(\omega)^{*}y_{j}v_{i_{j}}(\omega)$, so that $x_{j}\in\{a\}^{\prime}\cap{\mathcal{Q}}$. Then for each $m$, we can write $x_{j}=[X_{j}^{(n)}]_{n\in{\mathbb{N}}}$ for some sequence $X_{j}^{(n)}\in{\mathcal{D}}_{1/m}^{(n)}$ by Corollary 2.6. Then (2.9) implies that
$\left|\operatorname{tr}^{{\mathcal{Q}}}\left[v_{i_{1}}(\omega)x_{1}v_{i_{1}}(\omega)^{*}\dots v_{i_{k}}(\omega)x_{k}v_{i_{k}}(\omega))^{*}\right]\right|\leq 4k/m+2k\sqrt{24k\log(3Rm)/m}.$
Since $m$ was arbitrary, we obtain that $\operatorname{tr}^{{\mathcal{Q}}}[y_{1}\dots y_{m}]=0$, and so we have shown free independence of $\{u_{i}(\omega)\}^{\prime}$ for $i\in{\mathbb{N}}$. ∎
For the case of Theorem A, we present here a slightly more general version with each $B_{j}^{(n)}$ replaced by a tuple.
Theorem 2.12 (Asymptotic freeness of approximate commutants).
Let $U_{1}^{(n)}$, $U_{2}^{(n)}$, …be independent $n\times n$ Haar random unitary matrices. For each $j$, let $\mathbf{B}_{j}^{(n)}=(B_{j,\ell}^{(n)})_{\ell\in{\mathbb{N}}}$ be random matrices such that almost surely we have
$\limsup_{n\to\infty}\lVert B_{j,\ell}^{(n)}\rVert<\infty$
and
$\lim_{n\to\infty}\lVert[B_{j,\ell}^{(n)},U_{j}^{(n)}]\rVert_{2}=0.$
Then almost surely the tuples $\mathbf{B}_{j}^{(n)}$ are asymptotically freely independent, that is, for each alternating word $i_{1}\neq i_{2}\neq\dots\neq i_{k}$ and any non-commutative polynomials $p_{1}$, …, $p_{k}$, we have almost surely
(2.11) $\lim_{n\to\infty}\operatorname{tr}_{n}[(p_{1}(\mathbf{B}_{i_{1}}^{(n)})-\operatorname{tr}_{n}[p_{1}(\mathbf{B}_{i_{1}}^{(n)})])\dots(p_{k}(\mathbf{B}_{i_{k}}^{(n)})-\operatorname{tr}_{n}[p_{k}(\mathbf{B}_{i_{k}}^{(n)})])]=0.$
Proof.
Again, let $A^{(n)}$, $U_{j}^{(n)}$, and $V_{j}^{(n)}$ be as in Proposition 2.2.
Fix the word $i_{1}$, …, $i_{k}$. Let $C_{j}^{(n)}=p_{j}(\mathbf{B}_{i_{j}}^{(n)})$. Observe that almost surely $\limsup_{n\to\infty}\lVert C_{j}^{(n)}\rVert<\infty$. Moreover, since $\limsup_{n\to\infty}\lVert B_{j,k}^{(n)}\rVert<\infty$, we see that
$\limsup_{n\to\infty}\lVert[C_{j}^{(n)},U_{i_{j}}^{(n)}]\rVert_{2}=\limsup_{n\to\infty}\lVert[p_{j}(\mathbf{B}_{i_{j}}^{(n)}),U_{i_{j}}^{(n)}]\rVert_{2}=0,$
which follows from checking the case when $p_{j}$ is a monomial, which in turn follows from the triangle inequality and non-commutative Hölder inequality.
Let $D_{j}^{(n)}=(V_{i_{j}}^{(n)})^{*}C_{j}^{(n)}V_{i_{j}}^{(n)}$, and note that almost surely $\lVert[D_{j}^{(n)},A^{(n)}]\rVert_{2}\to 0$. Fix $m\in{\mathbb{N}}$. By Lemma 2.5, there is a matrix $D_{j,\epsilon}^{(n)}\in{\mathcal{D}}_{1/m}^{(n)}$ with
$\lVert D_{j,\epsilon}^{(n)}\rVert\leq 3\lVert D_{j}^{(n)}\rVert,\qquad\lVert D_{j,\epsilon}^{(n)}-D_{j}^{(n)}\rVert_{2}\leq m\lVert[D_{j}^{(n)},A^{(n)}]\rVert_{2},$
and it is obvious from the proof of that lemma that $D_{j,\epsilon}^{(n)}$ is a measurable function on the probability space. Recall that (2.9) holds almost surely by Lemma 2.10 and 2.11. Hence, almost surely
$\limsup_{n\to\infty}\left|\operatorname{tr}_{n}\left[V_{i_{1}}^{(n)}(D_{1,\epsilon}^{(n)}-\operatorname{tr}_{n}(D_{1,\epsilon}^{(n)}))(V_{i_{1}}^{(n)})^{*}\dots V_{i_{k}}^{(n)}(D_{k,\epsilon}^{(n)}-\operatorname{tr}_{n}(D_{k,\epsilon}^{(n)}))(V_{i_{k}}^{(n)})^{*}\right]\right|\\ \leq 4k/m+2k\sqrt{24k\log(3Rm)/m}.$
Since $\lVert D_{j,\epsilon}^{(n)}-D_{j}^{(n)}\rVert_{2}\to 0$ almost surely and also $\limsup_{n\to\infty}\lVert D_{j,\epsilon}^{(n)}\rVert<\infty$ almost surely, we obtain also that
$\limsup_{n\to\infty}\left|\operatorname{tr}_{n}\left[V_{i_{1}}^{(n)}(D_{1}^{(n)}-\operatorname{tr}_{n}(D_{1}^{(n)}))(V_{i_{1}}^{(n)})^{*}\dots V_{i_{k}}^{(n)}(D_{k}^{(n)}-\operatorname{tr}_{n}(D_{k}^{(n)}))(V_{i_{k}}^{(n)})^{*}\right]\right|\\ \leq 4k/m+2k\sqrt{24k\log(3Rm)/m}.$
Then since $m$ was arbitrary and since $p_{j}(\mathbf{B}_{i_{j}}^{(n)})=C_{j}^{(n)}=V_{i_{j}}^{(n)}D_{j}^{(n)}(V_{i_{j}}^{(n)})^{*}$, we obtain (2.11). ∎
3. Tools for the general approach
Toward the proof of Theorems C and F, we recall some results about model theory of operator algebras, as well as the version of Jung-Hayes $1$-bounded entropy [37, 27] for types developed in [34]. The reason that we use full types in the proof of our main theorem and not just existential types (which would correspond to the entropy in the presence of Hayes) is explained in Remark 4.3.
3.1. Model theory background
In the proof of Theorem C and Theorem F, we will use several concepts from model theory of tracial von Neumann algebras, in particular formulas, definable predicates, types, elementary submodels, and countable saturation. We explain below the minimal background for these concepts in tracial von Neumann algebras. For more general background on model theory for metric structures and tracial von Neumann algebras in particular, see [5] [26] [16], [23], [22], [34, §2-3], [36, §2].
Model theory of metric structures: In model theory of metric structures [5], a certain category of objects is formalized through a language ${\mathcal{L}}$ which describes the operations (functions from ${\mathcal{M}}^{n}$ to ${\mathcal{M}}$ such as addition or multiplication) and predicates (functions from ${\mathcal{M}}$ to ${\mathbb{R}}$ such as the trace or the distance), which can then be used to state axioms for the structures of interest. To state such axioms, one first defines formulas as expressions in formal variables $x_{1}$, $x_{2}$, …built from the operations and predicates in the language together with connectives and quantifiers, as explained in more detail below. Formulas with no free variables are called sentences, and a collection of sentences is called a theory.
Farah, Hart, and Sherman [17] described the language ${\mathcal{L}}_{\operatorname{tr}}$ for tracial von Neumann algebras, and a certain theory $\mathrm{T}_{\operatorname{tr}}$ that encodes the axioms of a tracial von Neumann algebra. In general, ${\mathcal{L}}$-structures are metric spaces equipped with functions corresponding to the operations and predicates in ${\mathcal{L}}$ (but which do not a priori satisfy any particular list of axioms, or theory). Thus, ${\mathcal{L}}_{\operatorname{tr}}$-structures have formal operations of addition, multiplication, *, and trace, but do not necessarily satisfy the *-algebra axioms. The ${\mathcal{L}}_{\operatorname{tr}}$-structures that are actually tracial von Neumann algebras are precisely those which satisfy $\mathrm{T}_{\operatorname{tr}}$.
Formulas: In continuous logic, the connectives are given by continuous functions. The quantifiers are supremum and infimum over appropriate sets called sorts or domains; the domains for ${\mathcal{L}}_{\operatorname{tr}}$ are the operator norm balls $B_{r}^{{\mathcal{M}}}$. ${\mathcal{L}}_{\operatorname{tr}}$-formulas in variables $\mathbf{x}=(x_{j})_{j\in{\mathbb{N}}}$ are formal expressions built up recursively as follows:
  • Basic formulas: A basic formula is an expression of the form $\operatorname{Re}\operatorname{tr}(t(\mathbf{x}))$ or $\lVert t(\mathbf{x})\rVert_{2}$, where $t$ is an expression formed through addition, multiplication, and *-operations (in practice, when evaluated on a von Neumann algebra, $t$ reduces to a *-polynomial  [^2] The definition of formula necessarily precedes writing axioms for von Neumann algebras. Thus, in this definition we don’t assume the *-algebra axioms, and hence, for instance, $x_{1}(x_{2}+x_{3})$ and $x_{1}x_{2}+x_{1}x_{3}$ are distinct expressions.).
  • Connectives: If $\phi_{1}(\mathbf{x})$, …, $\phi_{k}(\mathbf{x})$ are formulas, and $F:{\mathbb{R}}^{k}\to{\mathbb{R}}$ is a continuous function or connective, then $F(\phi_{1}(\mathbf{x}),\dots,\phi_{k}(\mathbf{x}))$ is a formula.
  • Quantifiers: Let $\phi(\mathbf{x},y)$ be a formula in variables $\mathbf{x}$ and another variable $y$. Fix $r>0$ and recall $B_{r}$ denotes the operator-norm ball of radius $r$. Then
    $\sup_{y\in B_{r}}\phi(\mathbf{x},y)\text{ and }\inf_{y\in B_{r}}\phi(\mathbf{x},y)$
    are formulas.
Given a formula $\phi$, a tracial von Neumann algebra ${\mathcal{M}}$, and $x_{1}$, $x_{2},\dots\in{\mathcal{M}}$, we can evaluate $\phi^{{\mathcal{M}}}(\mathbf{x})$ by substituting the actual elements $x_{j}$ instead of the formal variables, and evaluating each $\sup_{y\in B_{r}}$ or $\inf_{y\in B_{r}}$ symbol in the formula as the supremum or infimum over the ball $B_{r}^{{\mathcal{M}}}$ in ${\mathcal{M}}$. The mapping ϕ::superscriptitalic-$\phi^{{\mathcal{M}}}:{\mathcal{M}}^{{\mathbb{N}}}\to{\mathbb{R}}$ is called the interpretation of the formula.
An example of a formula for tracial von Neumann algebras is
$\phi(x)=\sup_{y\in B_{1}}\operatorname{tr}((xy-yx)^{*}(xy-yx)).$
For a tracial von Neumann algebra ${\mathcal{M}}$ and $x\in{\mathcal{M}}$, we have $\phi^{{\mathcal{M}}}(x)=0$ if and only if $x$ is in the center of ${\mathcal{M}}$.
Definable predicates: The set of formulas $\mathcal{F}$ forms an algebra over ${\mathbb{R}}$ because the formulas can be added and multiplied (the addition and multiplication functions ${\mathbb{R}}^{2}\to{\mathbb{R}}$ count as connectives). From a functional-analytic point of view, it is natural to complete the space of formulas into a Banach algebra (or something similar); see [26, §7.2]. In the case of ${\mathcal{L}}_{\operatorname{tr}}$, we can obtain a ‘‘Fréchet algebra’’ by taking the completion with respect to uniform convergence on each operator norm ball. More precisely, for a formula $\phi$ in variables $\mathbf{x}=(x_{j})_{j\in{\mathbb{N}}}$, for $\mathbf{r}=(r_{j})_{j\in{\mathbb{N}}}$, let
$\lVert\phi\rVert_{\mathbf{r}}=\sup\left\{|\phi^{{\mathcal{M}}}(\mathbf{x})|:{\mathcal{M}}\text{ tracial von Neumann algebra, }\mathbf{x}\in\prod_{j\in{\mathbb{N}}}B_{r_{j}}^{{\mathcal{M}}}\right\}.$
This defines a collection of seminorms, and the elements of the completion are called definable predicates relative to $\mathrm{T}_{\operatorname{tr}}$ (the notation here mentions $\mathrm{T}_{\operatorname{tr}}$ because we took the supremum only over ${\mathcal{M}}$ which satisfy $\mathrm{T}_{\operatorname{tr}}$ rather than all ${\mathcal{L}}_{\operatorname{tr}}$-structures). We will later use the fact that definable predicates satisfy a certain uniform continuity property with respect to $2$-norm.
Lemma 3.1 (See [34, Observation 3.11]).
Let $\phi$ be a definable predicate over $\mathcal{L}_{\operatorname{tr}}$ relative to $\mathrm{T}_{\operatorname{tr}}$ in countably many variables. Let $\mathbf{r}\in(0,\infty)^{{\mathbb{N}}}$ and $\epsilon>0$. Then there exists a finite $F\subseteq{\mathbb{N}}$ and $\delta>0$ such that, for every $\mathcal{M}\models\mathrm{T}_{\operatorname{tr}}$ and $\mathbf{x},\mathbf{y}\in\prod_{j\in{\mathbb{N}}}B_{r_{j}}^{{\mathcal{M}}}$,
$d^{{\mathcal{M}}}(x_{j},y_{j})<\delta\text{ for all }j\in F\implies|\phi^{{\mathcal{M}}}(\mathbf{x})-\phi^{{\mathcal{M}}}(\mathbf{y})|<\epsilon.$
Types: Naturally, we want to identify the algebra of definable predicates with the algebra of continuous functions on its Gelfand spectrum. This Gelfand spectrum is precisely the space of types. For each tuple $\mathbf{x}$, the type of $\mathbf{x}$ is the mapping $\operatorname{tp}^{{\mathcal{M}}}(\mathbf{x}):\phi\mapsto\phi^{{\mathcal{M}}}(\mathbf{x})$ from the algebra of formulas (and more generally definable predicates) to the real numbers. We denote the set of types of countable tuples in tracial von Neumann algebras by $\mathbb{S}(\mathrm{T}_{\operatorname{tr}})$. For $\mathbf{r}\in(0,\infty)^{{\mathbb{N}}}$, the set of types of tuples $\mathbf{x}$ with $x_{j}\in B_{r_{j}}$ will be denoted $\mathbb{S}_{\mathbf{r}}(\mathrm{T}_{\operatorname{tr}})$. We equip $\mathbb{S}_{\mathbf{r}}(\mathrm{T}_{\operatorname{tr}})$ with the weak-* topology (also known as the logic topology), and we equip $\mathbb{S}(\mathrm{T})$ with the inductive limit topology obtained by viewing it as the union of the spaces $\mathbb{S}_{\mathbf{r}}(\mathrm{T}_{\operatorname{tr}})$ (see [34, §3.1]). With these definitions in hand, the algebra of definable predicates is isomorphic to $C(\mathbb{S}(\mathrm{T}_{\operatorname{tr}}))$ and the norm $\lVert\cdot\rVert_{\mathbf{r}}$ given above coincides with the $C(\mathbb{S}_{\mathbf{r}}(\mathrm{T}_{\operatorname{tr}}))$ norm.
The space $\mathbb{S}_{\mathbf{r}}(\mathrm{T}_{\operatorname{tr}})$ is compact Hausdorff and metrizable [34, Observation 3.14]. In particular, for any point $\mu\in\mathbb{S}_{\mathbf{r}}(\mathrm{T}_{\operatorname{tr}})$, there is a nonnegative continuous function $f$ that vanishes only at $\mu$. By [36, Lemma 2.16], $f$ automatically extends to a continuous function on $\mathbb{S}(\mathrm{T}_{\operatorname{tr}})$, that is, a definable predicate. We record this result for later use.
Lemma 3.2.
Let $\mu\in\mathbb{S}_{\mathbf{r}}(\mathrm{T}_{\operatorname{tr}})$ be a type. Then there exists a nonnegative definable predicate $\phi$ such that whenever ${\mathcal{M}}$ is a tracial von Neumann algebra and $\mathbf{x}\in\prod_{j\in{\mathbb{N}}}B_{r_{j}}^{{\mathcal{M}}}$, we have
$\phi^{{\mathcal{M}}}(\mathbf{x})=0\iff\operatorname{tp}^{{\mathcal{M}}}(\mathbf{x})=\mu.$
In the proof of Theorem C, we will be concerned with the limiting type of a countable tuple of independent Haar random unitary matrices, as well as the limiting types of deterministic matrices. We first recall Łoś’s theorem in the setting of continuous logic for tracial von Neumann algebras.
Theorem 3.3 (See [5, Theorem 5.4]).
Let $({\mathcal{M}}_{i})_{i\in I}$ be tracial von Neumann algebras indexed by a set $I$. Let ${\mathcal{U}}$ be a non-principal ultrafilter on $I$. Let $\mathbf{r}\in(0,\infty)^{{\mathbb{N}}}$, and for each $i,j\in I$, let $x_{i,j}\in B_{r_{j}}^{{\mathcal{M}}_{i}}$. Let $\mathbf{x}_{i}=(x_{i,j})_{j\in{\mathbb{N}}}$. Let ${\mathcal{M}}=\prod_{i\to{\mathcal{U}}}{\mathcal{M}}_{i}$ be the ultraproduct, and let $y_{j}$ be the equivalence class $[x_{i,j}]_{i\in I}$ in ${\mathcal{M}}$, and let $\mathbf{y}=(y_{j})_{j\in{\mathbb{N}}}$. Then for every formula $\phi$ in countably many variables, we have
$\phi^{{\mathcal{M}}}(\mathbf{y})=\lim_{i\to{\mathcal{U}}}\phi^{{\mathcal{M}}_{i}}(\mathbf{x}_{i}).$
Hence also $\lim_{i\to{\mathcal{U}}}\operatorname{tp}^{{\mathcal{M}}_{i}}(\mathbf{x}_{i})=\operatorname{tp}^{{\mathcal{M}}}(\mathbf{y})$.
We will also rely on the fact that the limit of the type of a tuple of Haar random unitary matrices exists almost surely, which follows from concentration of measure. The proof is the same as [18, Lemma 4.2], so we only give a sketch below. In the following, we denote the dual pairing of a type $\mu$ with a formula $\phi$ by $\mu[\phi]$.
Lemma 3.4.
Fix a free ultrafilter ${\mathcal{U}}$ on ${\mathbb{N}}$. Let $U_{1}^{(n)}$, $U_{2}^{(n)}$, …be independent $n\times n$ Haar random unitaries. For each ${\mathcal{L}}_{\operatorname{tr}}$-formula $\phi$ in countably many variables, let
$\mu_{\operatorname{Haar}}[\phi]=\lim_{n\to{\mathcal{U}}}\mathbb{E}\phi^{M_{n}({\mathbb{C}})}(U_{1}^{(n)},U_{2}^{(n)},\dots).$
Then almost surely $\lim_{n\to{\mathcal{U}}}\phi^{M_{n}({\mathbb{C}})}(U_{1}^{(n)},U_{2}^{(n)},\dots)=\mu_{\operatorname{Haar}}[\phi]$. In other words,
$\operatorname{tp}^{M_{n}({\mathbb{C}})}(U_{1}^{(n)},U_{2}^{(n)},\dots)\to\mu_{\operatorname{Haar}}$
weak-* as $n\to{\mathcal{U}}$ almost surely.
Sketch of proof.
Let $\mathbf{r}=(1,1,\dots)$. Note that every ${\mathcal{L}}_{\operatorname{tr}}$-formula $\phi$ can be approximated in $\lVert\cdot\rVert_{\mathbf{r}}$ by one which is uniformly Lipschitz (for all tracial von Neumann algebras and all inputs); this follows by taking the connectives $F:{\mathbb{R}}^{k}\to{\mathbb{R}}$ to be Lipschitz; see [18, proof of Lemma 4.2]. By a ‘‘$3\epsilon$ argument’’ it suffices to check the claim when $\phi$ is Lipschitz. In this case, using concentration of measure (Lemma 2.8) and the Borel-Cantelli lemma, we see that almost surely
$\lim_{n\to\infty}|\phi^{M_{n}({\mathbb{C}})}(U_{1}^{(n)},U_{2}^{(n)},\dots)-\mathbb{E}\phi^{M_{n}({\mathbb{C}})}(U_{1}^{(n)},U_{2}^{(n)},\dots)|=0,$
which implies the claim of the lemma. ∎
Remark 3.5.
Note that $\mu_{\operatorname{Haar}}$ may depend on the choice of ultrafilter ${\mathcal{U}}$, since we do not even know whether the matrix ultraproducts for different ultrafilters are elementarily equivalent. Very little is known at this point about the large $n$ behavior of formulas containing quantifiers on Haar unitaries. See [35, §5.2] and [18, §4] for related discussion and results.
Elementary substructures: If ${\mathcal{L}}$ is a metric language, and ${\mathcal{M}}$ and ${\mathcal{N}}$ are ${\mathcal{L}}$-structures, we say that ${\mathcal{M}}$ is an elementary submodel of ${\mathcal{N}}$ if for every formula $\phi$ and for every tuple $\mathbf{x}$ in ${\mathcal{M}}$, we have that $\phi^{{\mathcal{M}}}(\mathbf{x})=\phi^{{\mathcal{N}}}(\mathbf{x})$. Equivalently, $\operatorname{tp}^{{\mathcal{M}}}(\mathbf{x})=\operatorname{tp}^{{\mathcal{N}}}(\mathbf{x})$. We will use the following fact, known as the Downward Löwenheim-Skolem theorem; see [5, Proposition 7.3]. We state it for convenience in the particular case of tracial von Neumann algebras.
Proposition 3.6 (Downward Löwenheim-Skolem theorem).
Let ${\mathcal{M}}$ be a tracial von Neumann algebra, and let ${\mathcal{A}}\subseteq{\mathcal{M}}$ be a separable von Neumann subalgebra. Then there exists a separable elementary substructure $\widehat{{\mathcal{M}}}\preceq{\mathcal{M}}$ that contains ${\mathcal{A}}$.
Countable saturation: Countable saturation of an ${\mathcal{L}}$-structure means essentially that sets of formulas in countably many variables that admit approximate solutions must admit exact solutions.
Let $\Phi$ be a set of ${\mathcal{L}}_{\operatorname{tr}}$-formulas in countably many variables $(x_{j})_{j\in{\mathbb{N}}}$ and parameters or constants $(a_{j})_{j\in{\mathbb{N}}}$ from some tracial von Neumann algebra ${\mathcal{M}}$. Fix some $\mathbf{r}\in(0,\infty)^{{\mathbb{N}}}$.
  • We say that $\Phi$ is satisfiable in $\prod_{j\in{\mathbb{N}}}B_{r_{j}}^{{\mathcal{M}}}$ if there exists $\mathbf{x}\in\prod_{j\in{\mathbb{N}}}B_{r_{j}}^{{\mathcal{M}}}$ such that $\phi^{{\mathcal{M}}}(\mathbf{x},\mathbf{a})=0$ for all $\phi\in\Phi$.
  • We say that $\Phi$ is finitely approximately satisfiable in $\prod_{j\in{\mathbb{N}}}B_{r_{j}}^{{\mathcal{M}}}$ if for every $\phi_{1}$, …, $\phi_{k}$ in $\Phi$ and every $\epsilon$, there exists some $\mathbf{x}\in\prod_{j\in{\mathbb{N}}}B_{r_{j}}^{{\mathcal{M}}}$ satisfying $|\phi_{i}(\mathbf{x},\mathbf{a})|<\epsilon$ for $i=1$, …, $k$.
  • We say that ${\mathcal{M}}$ is countably saturated (or $\aleph_{1}$-saturated) if for every $\mathbf{r}$, and for every set of formulas $\Phi$ in countably many variables and countably many parameters $\mathbf{a}$, if $\Phi$ is finitely approximately satisfiable in $\prod_{j\in{\mathbb{N}}}B_{r_{j}}^{{\mathcal{M}}}$, then $\Phi$ is satisfiable in $\prod_{j\in{\mathbb{N}}}B_{r_{j}}^{{\mathcal{M}}}$.
Most ultraproducts are countably saturated. Recall that an ultrafilter ${\mathcal{U}}$ on a set $I$ is said to be countably incomplete if there is a countable family of sets in ${\mathcal{U}}$ with empty intersection. We recall the following fact:
Lemma 3.7 (See [16, Proposition 4.11], [5, Proposition 7.6]).
Let ${\mathcal{M}}=\prod_{i\to{\mathcal{U}}}{\mathcal{M}}_{i}$ for some tracial von Neumann algebras $({\mathcal{M}}_{i})_{i\in I}$ where $I$ is an infinite index set and ${\mathcal{U}}$ is a countably incomplete ultrafilter on $I$. Then ${\mathcal{M}}$ is countably saturated.
3.2. $1$-bounded entropy for types
In free entropy theory, $1$-bounded entropy is a notion of metric entropy for matricial approximations defined by Hayes [27] and inspired by the work of Jung [37]. Here we describe the version for full types from [34].
If ${\mathcal{O}}$ is a subset of the type space $\mathbb{S}(\mathrm{T}_{\operatorname{tr}})$ and $\mathbf{r}\in(0,\infty)^{{\mathbb{N}}}$, we define
$\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}})=\left\{\mathbf{X}\in\prod_{j\in{\mathbb{N}}}D_{r_{j}}^{M_{n}({\mathbb{C}})}:\operatorname{tp}^{M_{n}({\mathbb{C}})}(\mathbf{X})\in{\mathcal{O}}\right\}.$
We view this as a microstate space as in Voiculescu’s free entropy theory [61]. Entropy of types is defined by the exponential growth rates of covering numbers of these spaces $\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}})$ up to unitary conjugation.
Definition 3.8 (Orbital covering numbers).
Given $\Omega\subseteq M_{n}({\mathbb{C}})^{{\mathbb{N}}}$ and a finite $F\subseteq{\mathbb{N}}$ and $\epsilon>0$, we define $N_{F,\epsilon}^{\operatorname{orb}}(\Omega)$ to be the set of $\mathbf{Y}\in M_{n}({\mathbb{C}})^{{\mathbb{N}}}$ such that there exists a unitary $U$ in $M_{n}({\mathbb{C}})$ and $\mathbf{X}\in\Omega$ such that $\lVert Y_{i}-UX_{i}U^{*}\rVert_{2}<\epsilon$ for all $i\in F$. If $\Omega\subseteq N_{F,\epsilon}^{\operatorname{orb}}(\Omega^{\prime})$, we say that $\Omega^{\prime}$ orbitally $(F,\epsilon)$-covers $\Omega$. We denote by $K_{F,\epsilon}^{\operatorname{orb}}(\Omega)$ the minimum cardinality of a set $\Omega^{\prime}$ that orbitally $(F,\epsilon)$-covers $\Omega$.
Definition 3.9 ($1$-bounded entropy for types [34]).
Fix a non-principal ultrafilter ${\mathcal{U}}$ on ${\mathbb{N}}$. For $\mu\in\mathbb{S}(\mathrm{T}_{\operatorname{tr}})$ and $F\subseteq I$ finite and $\epsilon>0$, we define
$\operatorname{Ent}_{\mathbf{r},F,\epsilon}^{\mathcal{U}}(\mu)=\inf_{\text{open}{\mathcal{O}}\ni\mu}\lim_{n\to{\mathcal{U}}}\frac{1}{n^{2}}\log K_{F,\epsilon}^{\operatorname{orb}}(\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}})).$
Then let
$\operatorname{Ent}_{\mathbf{r}}^{{\mathcal{U}}}(\mu):=\sup_{\begin{subarray}{c}\text{finite }F\subseteq{\mathbb{N}}\\ \epsilon>0\end{subarray}}\operatorname{Ent}_{\mathbf{r},F,\epsilon}^{{\mathcal{U}}}(\mu).$
and
$\operatorname{Ent}^{{\mathcal{U}}}(\mu):=\sup_{\mathbf{r}\in(0,\infty)^{{\mathbb{N}}}}\operatorname{Ent}_{\mathbf{r}}^{{\mathcal{U}}}(\mu).$
We also remark that the same definitions make sense for types of finite tuples $(x_{1},\dots,x_{m})$ instead of countable tuples. In that setting, one does not need to consider the finite subset $F$ since $F$ can be taken to be $\{1,\dots,m\}$.
By [34, Corollary 4.10], if $\mathbf{x}$ and $\mathbf{y}$ are tuples in a tracial von Neumann algebra ${\mathcal{M}}$ and $\mathrm{W}^{*}(\mathbf{x})=\mathrm{W}^{*}(\mathbf{y})$, then $\operatorname{tp}^{{\mathcal{M}}}(\mathbf{x})$ and $\operatorname{tp}^{{\mathcal{M}}}(\mathbf{y})$ have the same entropy. Hence, for separable ${\mathcal{A}}\subseteq{\mathcal{M}}$, one can define $\operatorname{Ent}^{{\mathcal{U}}}({\mathcal{A}}:{\mathcal{M}})$ as the entropy of any generating tuple for ${\mathcal{A}}$. More generally, if ${\mathcal{A}}$ is not necessarily separable, one defines
$\operatorname{Ent}^{{\mathcal{U}}}({\mathcal{A}}:{\mathcal{M}}):=\sup\{\operatorname{Ent}^{{\mathcal{U}}}(\operatorname{tp}^{{\mathcal{M}}}(\mathbf{x})):\mathbf{x}\in{\mathcal{A}}^{{\mathbb{N}}}\},$
and by [34, Observation 4.12] this agrees with the entropy of any countable (or finite) generating tuple in the case when ${\mathcal{A}}$ is separable. Now if ${\mathcal{A}}\subseteq{\mathcal{B}}\subseteq{\mathcal{M}}$, then
(3.1) $\operatorname{Ent}^{{\mathcal{U}}}({\mathcal{A}}:{\mathcal{M}})\leq\operatorname{Ent}^{{\mathcal{U}}}({\mathcal{B}}:{\mathcal{M}}).$
Entropy for types versus entropy in the presence: The relationship between the entropy for types described above and Hayes’ $1$-bounded entropy is as follows (see [34, §5] for details). Hayes’ $1$-bounded entropy $h^{{\mathcal{U}}}({\mathcal{A}}:{\mathcal{M}})$ of ${\mathcal{A}}$ in the presence of ${\mathcal{M}}$ (with respect to the ultrafilter ${\mathcal{U}}$) arises by looking at microstate spaces for the existential type of tuples $\mathbf{x}$ rather than the full type. The existential type describes the evaluation on $\mathbf{x}$ of formulas of the form $\phi(\mathbf{x})=\inf_{y\in B_{r}}\psi(\mathbf{x},y)$, where $\psi$ is a quantifier-free formula (i.e. a formula with no supremum or infimum in it). The space of existential types is equipped with a topology that is non-Hausdorff because a basic neighborhood is defined by one-sided upper bounds on a finite family of inf-formulas. The microstate spaces defined by a neighborhood of the existential type are a special case of microstate spaces defined by neighborhoods of the full type, and hence
(3.2) $\operatorname{Ent}^{{\mathcal{U}}}({\mathcal{A}}:{\mathcal{M}})\leq h^{{\mathcal{U}}}({\mathcal{A}}:{\mathcal{M}});$
see e.g. [34, Lemma 5.13].
Orbital versus relative entropy: There are actually two approaches to defining metric entropy–one based on covering numbers up to unitary conjugation, and one based on covering number of microstate spaces relative to a fixed microstate sequence $A^{(n)}$ for a normal element $a$ with diffuse spectrum. These were shown to be equivalent in [27, Lemma A.5] for the setting of $a$ self-adjoint and for the $1$-bounded entropy in the presence $h$. We need the analogous result for the entropy for types, and with using a Haar unitary instead of a self-adjoint element for the fixed microstate. We include a self-contained proof for convenience; the approach here is slightly different since we fix a very specific $A^{(n)}$ and use Lemma 2.5 rather than using Szarek’s covering estimates for Grassmannians [59].
Lemma 3.10.
Let $\mu\in\mathbb{S}_{\mathbf{r}}(\mathrm{T}_{\operatorname{tr}})$ be the type of some infinite tuple $(a,x_{1},x_{2},\dots)$ such that $a$ is a Haar unitary. Let $A^{(n)}=\operatorname{diag}(1,\zeta_{n},\dots,\zeta_{n}^{n-1})$, where $\zeta_{n}=e^{2\pi i/n}$. For a neighborhood ${\mathcal{O}}$ of $\mu$, let
$\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}};A^{(n)}\rightsquigarrow a)=\{\mathbf{X}^{(n)}\in\prod_{j\in{\mathbb{N}}}D_{r_{j}}^{M_{n}({\mathbb{C}})}:\operatorname{tp}^{M_{n}({\mathbb{C}})}(A^{(n)},\mathbf{X}^{(n)})\in{\mathcal{O}}\}.$
Then
$\operatorname{Ent}_{\mathbf{r}}^{{\mathcal{U}}}(\mu)=\sup_{(F,\epsilon)}\inf_{{\mathcal{O}}\ni\mu}\lim_{n\to{\mathcal{U}}}\frac{1}{n^{2}}\log K_{\epsilon}(\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}};A^{(n)}\rightsquigarrow a)).$
Proof.
First, let us show $\leq$. Given $(F,\epsilon)$ and given a neighborhood ${\mathcal{O}}$ of $\mu$, we claim that there is a neighborhood ${\mathcal{O}}^{\prime}$ such that $\Gamma_{(1,\mathbf{r})}({\mathcal{O}}^{\prime})$ is contained in the orbital $(F,\epsilon/2)$-neighborhood of $\{A^{(n)}\}\times\Gamma_{\mathbf{r}}({\mathcal{O}};A^{(n)}\rightsquigarrow a)$.
By Lemma 3.2, there exists a definable predicate $\phi$ such that, when ${\mathcal{M}}$ is a tracial von Neumann algebra and $\mathbf{y}\in\prod_{j\in{\mathbb{N}}}B_{r_{j}}^{{\mathcal{M}}}$, then $\phi^{{\mathcal{M}}}(b,\mathbf{y})=0$ if and only if $\operatorname{tp}^{{\mathcal{M}}}(b,\mathbf{y})=\mu$. Note that there exists $m\in{\mathbb{N}}$ such that if $\phi^{{\mathcal{M}}}(b,\mathbf{y})\leq 1/m$, then $\operatorname{tp}^{{\mathcal{M}}}(\mathbf{y})\in{\mathcal{O}}$. Indeed, let $\mathcal{K}_{m}$ be the set of types in $\mathbb{S}_{\mathbf{r}}(\mathrm{T}_{\operatorname{tr}})\setminus{\mathcal{O}}$ that satisfy $\varphi\leq 1/m$, which is a closed subset of $\mathbb{S}_{\mathbf{r}}(\mathrm{T}_{\operatorname{tr}})$. If $\mathcal{K}_{m}$ were nonempty for all $m$, then by compactness $\bigcap_{m\in{\mathbb{N}}}\mathcal{K}_{m}$ would be nonempty, which would yield some ${\mathcal{M}}$ and $(b,\mathbf{y})$ such that $\phi^{{\mathcal{M}}}(b,\mathbf{y})=0$ and $\operatorname{tp}^{{\mathcal{M}}}(b,\mathbf{y})\not\in{\mathcal{O}}$, which contradicts the choice of $\phi$. Hence, we have that $\phi^{{\mathcal{M}}}(b,\mathbf{y})\leq 1/m$ implies $\operatorname{tp}^{{\mathcal{M}}}(b,\mathbf{y})\in{\mathcal{O}}$.
Using uniform continuity (Lemma 3.1), there exists $\delta>0$ such that for all tracial von Neumann algebras ${\mathcal{M}}$ and for $a,b\in B_{r_{1}}^{{\mathcal{M}}}$ and $x_{j}\in B_{r_{j}}^{{\mathcal{M}}}$, $j\geq 2$, we have
$\lVert a-b\rVert_{2}<\delta\implies|\phi^{{\mathcal{M}}}(a,\mathbf{x})-\phi^{{\mathcal{M}}}(b,\mathbf{x})|<\frac{1}{2m}.$
Let $k$ be such that $4\pi/k<\min(\delta,\epsilon/2)$, and let $\mathcal{O}_{k}$ be the neighborhood of the Haar measure on the unit circle described in Lemma 2.1; then since the weak-* topology on $\mathcal{P}(S^{1})$ is given by testing against trigonometric polynomials, there exist *-polynomials $p_{1}$, …, $p_{\ell}$ such that
  • $\operatorname{Re}\operatorname{tr}(p_{j}(a))<0$ for all $j$;
  • for any unitary $u$, if $\operatorname{Re}\operatorname{tr}(p_{j}(u))<0$, for all $j$, then the spectral distribution of $u$ is in ${\mathcal{O}}_{k}$.
Now let ${\mathcal{O}}^{\prime}$ be the set of types $\operatorname{tp}^{{\mathcal{M}}}(b,\mathbf{y})$ satisfying
$\phi^{{\mathcal{M}}}(b,\mathbf{y})<\frac{1}{2m}\text{ and }\max_{j\in[\ell]}\operatorname{Re}\operatorname{tr}(p_{j}(b))<0.$
Suppose that $(B,\mathbf{Y})$ is an $n\times n$ matrix tuple whose type is in $\mathcal{O}^{\prime}$. Then the spectral measure of $B$ is in ${\mathcal{O}}_{k}$, and therefore by Lemma 2.1, there is a unitary $V$ such that $\lVert V^{*}BV-A^{(n)}\rVert_{2}<4\pi/k<\min(\delta,\epsilon/2)$. Let $\mathbf{X}=V\mathbf{Y}V^{*}$. Since $\lVert A^{(n)}-VBV^{*}\rVert_{2}<\delta$, we have
$\phi^{M_{n}({\mathbb{C}})}(A^{(n)},\mathbf{X})<\phi^{M_{n}({\mathbb{C}})}(VBV^{*},\mathbf{X})+\frac{1}{2m}=\phi^{M_{n}}(B,\mathbf{Y})+\frac{1}{2m}<\frac{1}{m},$
where we used the unitary invariance of the formula $\phi$ and the assumption that $\operatorname{tp}^{M_{n}({\mathbb{C}})}(B,\mathbf{Y})\in{\mathcal{O}}^{\prime}$. Therefore, $\operatorname{tp}^{M_{n}({\mathbb{C}})}(A^{(n)},\mathbf{X})\in{\mathcal{O}}$, and $(B,\mathbf{Y})$ is in the orbital $\epsilon/2$-neighborhood of $(A^{(n)},\mathbf{X})$. Thus, $\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}})$ is in the orbital $(F,\epsilon/2)$ neighborhood of $\{A^{(n)}\}\times\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}}\mid A^{(n)}\rightsquigarrow a)$. Hence,
$K_{\epsilon}^{\operatorname{orb}}(\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}}^{\prime}))\leq K_{\epsilon/2}(\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}};A^{(n)}\rightsquigarrow a)).$
From this we obtain
$\operatorname{Ent}_{\mathbf{r},F,\epsilon}^{{\mathcal{U}}}(\mu)\leq\inf_{{\mathcal{O}}\ni\mu}\lim_{n\to{\mathcal{U}}}\frac{1}{n^{2}}\log K_{\epsilon/2}(\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}};A^{(n)}\rightsquigarrow a)),$
and then taking the supremum over $(F,\epsilon)$ finishes the inequality $\leq$.
For the other direction, again fix $(F,\epsilon)$ and ${\mathcal{O}}$. We want to bound the $(F,\epsilon)$ covering number of $\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}}\mid A^{(n)}\rightsquigarrow a)$ in terms of the orbital $(F,\delta)$ covering number of $\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}})$ where $\delta\in(0,1)$ is to be chosen later. Let $\Omega$ be a set that orbitally $(F,\delta)$-covers $\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}})$. In particular, $\{A^{(n)}\}\times\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}}\mid A^{(n)}\rightsquigarrow a)$ is covered by the orbital neighborhoods $N_{(F,\delta)}^{\operatorname{orb}}(B,\mathbf{X})$ for $(B,\mathbf{X})\in\Omega$. Thus, our goal is to estimate the plain covering number of the sets
$S_{(B,\mathbf{X})}=[\{A^{(n)}\}\times\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}}\mid A^{(n)}\rightsquigarrow a)]\cap N_{(F,\delta)}^{\operatorname{orb}}(B,\mathbf{X}).$
The idea is that unitaries that we conjugate by must approximately fix $A^{(n)}$ and hence are approximately band matrices. More precisely, if $(A^{(n)},X_{1})$ and $(A^{(n)},X_{2})$ are both in $S_{(B,\mathbf{X})}$, then there is a unitary $V$ such that
$(A^{(n)},X_{1})\in N_{(F,\delta)}(VA^{(n)}V^{*},VX_{2}V^{*}).$
In particular, $\lVert[V,A^{(n)}]\rVert_{2}<\delta$, and so by Lemma 2.5,
$d(V,\mathcal{D}_{\sqrt{\delta}}^{(n)}\cap B_{3}^{M_{n}({\mathbb{C}})})\leq\frac{8\sqrt{\pi}}{\sqrt{\delta}}2\delta=16\sqrt{\pi\delta}.$
Thus, by Lemma 2.11,
$K_{\sqrt{\delta}+16\sqrt{\pi\delta}}(\{V\in\mathbb{U}_{n}:\lVert[V,A^{(n)}]\rVert_{2}<\delta\})\leq K_{\sqrt{\delta}}(\mathcal{D}_{\sqrt{\delta}}^{(n)}\cap B_{3}^{M_{n}({\mathbb{C}}))})\leq\exp(n^{2}\cdot 2\sqrt{\delta}\log(3R/\sqrt{\delta})).$
Hence, fixing some $(A^{(n)},X_{0})\in S_{(B,\mathbf{X})}$, we have that
$S_{(B,\mathbf{X})}\subseteq N_{(F,2\delta)}(\{(A^{(n)},VX_{0}V^{*}):V\in\mathbb{U}_{n}:\lVert[V,A^{(n)}]\rVert_{2}<\delta\}).$
Then since $\lVert(X_{0})_{j}\rVert\leq r_{j}$, we have that $V\mapsto V(X_{0})_{j}V^{*}$ is $2r_{j}$-Lipschitz. Let $r_{F}=\max_{j\in F}r_{j}$. Then any $(F,\sqrt{\delta}+16\sqrt{\pi\delta})$-covering of $\{V\in\mathbb{U}_{n}:\lVert[V,A^{(n)}]\rVert_{2}<\delta\}$ yields a $(F,r_{F}(\sqrt{\delta}+16\sqrt{\pi\delta}))$-covering of the conjugation orbit of $(A^{(n)},X_{0})$ and so a $(F,2\delta+r_{F}(\sqrt{\delta}+16\sqrt{\pi\delta}))$-covering of $S_{(B,\mathbf{X})}$. Hence, choose $\delta$ small enough that $2\delta+r_{F}(\sqrt{\delta}+16\sqrt{\pi\delta})<\epsilon$. Then we obtain
$K_{F,\epsilon}(\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}}\mid A^{(n)}\rightsquigarrow a))\leq\sum_{(B,\mathbf{X})\in\Omega}K_{F,\epsilon}(S_{(B,\mathbf{X})})\leq K_{F,\delta}^{\operatorname{orb}}(\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}}))\exp(n^{2}\cdot 2\sqrt{\delta}\log(3R/\sqrt{\delta})).$
Hence,
$\frac{1}{n^{2}}\log K_{F,\epsilon}(\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}}\mid A^{(n)}\rightsquigarrow a))\leq\frac{1}{n^{2}}\log K_{F,\delta}^{\operatorname{orb}}(\Gamma_{\mathbf{r}}^{(n)}({\mathcal{O}}))+2\sqrt{\delta}\log(3R/\sqrt{\delta}).$
Since $\delta$ can be chosen arbitarily small, we get the desired inequality. ∎
4. General freeness results
We are now ready to prove Theorem C and F.
4.1. Free independence phenomena in matrix ultraproducts
In this section, we fix a countably incomplete ultrafilter ${\mathcal{U}}$ on ${\mathbb{N}}$, and write ${\mathcal{Q}}=\prod_{n\to{\mathcal{U}}}M_{n}({\mathbb{C}})$.
Lemma 4.1.
Let $U_{1}^{(n)}$, $U_{2}^{(n)}$, …be independent Haar unitaries. Let $k\in{\mathbb{N}}$, and let $i_{1}\neq i_{2}\neq\dots\neq i_{k}$. For each $j=1$, …, $k$, let $\nu_{j}\in\mathbb{S}_{(1,1)}(\mathrm{T}_{\operatorname{tr}})$ be the type of some pair $(u,x)$ in ${\mathcal{Q}}$ where $u$ is a Haar unitary and $x$ has trace zero, such that $\operatorname{Ent}^{{\mathcal{U}}}(\nu_{j})=0$. Then almost surely, for all $x_{j}$ with $\operatorname{tp}^{{\mathcal{Q}}}(u_{i_{j}}(\omega),x_{j})=\nu_{j}$, we have $\operatorname{tr}^{{\mathcal{Q}}}(x_{1}\dots x_{k})=0$.
Proof.
By Proposition 2.2, let $U_{j}^{(n)}$ be an independent family of Haar random unitaries and $V_{j}^{(n)}$ another independent family of Haar random unitaries such that $\lVert U_{j}^{(n)}-V_{j}^{(n)}A^{(n)}(V_{j}^{(n)}\rVert)^{*}\to 0$ almost surely, where $A^{(n)}=\operatorname{diag}(1,\zeta_{n},\dots,\zeta_{n}^{n-1})$. For each outcome $\omega$, let $v_{j}(\omega)$ be the corresponding element of the ultraproduct ${\mathcal{Q}}$, so that $u_{j}(\omega)=v_{j}(\omega)av_{j}(\omega)^{*}$.
Fix $m\in{\mathbb{N}}$. Since $\operatorname{Ent}^{{\mathcal{U}}}(\nu_{j})=0$, using Lemma 3.10 for the case of a $1$-tuple rather than a countable tuple, there exists some neighborhood ${\mathcal{O}}_{j,m}$ of $\nu_{j}$ such that
$\lim_{n\to{\mathcal{U}}}\frac{1}{n^{2}}\log K_{1/m}(\Gamma_{1}^{(n)}({\mathcal{O}}_{j,m}\mid A^{(n)}\rightsquigarrow a))<\frac{1}{m}.$
(Here we take $\epsilon=1/m$ and $F=\{1\}$.) Then applying Lemma 2.10 with $\delta=\epsilon=1/m$, we see that almost surely
$\lim_{n\to{\mathcal{U}}}\sup_{X_{1}\in\Gamma_{1}^{(n)}({\mathcal{O}}_{1,m}\mid A^{(n)}\rightsquigarrow a)}\dots\sup_{X_{k}\in\Gamma_{1}^{(n)}({\mathcal{O}}_{k,m}\mid A^{(n)}\rightsquigarrow a)}\\ \left|\operatorname{tr}_{n}\left[V_{i_{1}}^{(n)}(X_{1}-\operatorname{tr}_{n}(X_{1}))(V_{i_{1}}^{(n)})^{*}\dots V_{i_{k}}^{(n)}(X_{k}-\operatorname{tr}_{n}(X_{k}))(V_{i_{k}}^{(n)})^{*}\right]\right|\\ \leq 4k/m+2k\sqrt{12k/m}.$
By taking the countable intersection, almost surely this holds for all $m\in{\mathbb{N}}$.
Now fix an outcome $\omega$ in the almost sure event where the above inequality holds for all $m$. Suppose that $\operatorname{tp}^{{\mathcal{Q}}}(u_{i_{j}}(\omega),y_{j})=\nu_{j}$. Write $y_{j}=v_{i_{j}}(\omega)x_{j}v_{i_{j}}(\omega)^{*}$, and represent $X_{j}=[X_{j}^{(n)}]_{n\in{\mathbb{N}}}$ with $\lVert X_{j}^{(n)}\rVert\leq 1$. By Łoś’s theorem (Theorem 3.3),
$\lim_{n\to{\mathcal{U}}}\operatorname{tp}^{M_{n}({\mathbb{C}})}(A^{(n)},x_{j}^{(n)})=\operatorname{tp}^{{\mathcal{Q}}}(a,x_{j})=\operatorname{tp}^{{\mathcal{Q}}}(u_{i_{j}}(\omega),y_{j})=\nu_{j}.$
Thus, there is some $A\in{\mathcal{U}}$ such that $\operatorname{tp}^{M_{n}({\mathbb{C}})}(A^{(n)},X_{j}^{(n)})\in{\mathcal{O}}_{m,j}$ for each $j=1$, …, $k$ for each $n\in A$. Hence,
$|\operatorname{tr}^{{\mathcal{Q}}}(y_{1}\dots y_{k})|=|\operatorname{tr}^{{\mathcal{Q}}}(v_{i_{1}}(\omega)x_{1}v_{i_{1}}(\omega)^{*}\dots v_{i_{k}}(\omega)x_{k}v_{i_{k}}(\omega)^{*})|\\ =\lim_{n\to{\mathcal{U}}}\left|\operatorname{tr}_{n}\left[V_{i_{1}}^{(n)}(\omega)(X_{1}-\operatorname{tr}_{n}(X_{1}))(V_{i_{1}}^{(n)}(\omega))^{*}\dots V_{i_{k}}^{(n)}(\omega)(X_{k}-\operatorname{tr}_{n}(X_{k}))(V_{i_{k}}^{(n)}(\omega))^{*}\right]\right|\\ \leq 4k/m+2k\sqrt{12k/m}.$
Then since $m$ was arbitrary, we get $\operatorname{tr}^{{\mathcal{Q}}}(y_{1}\dots y_{k})=0$ as desired. ∎
In the proof of Theorem 4.4 below, we will want to apply the conclusion of Lemma 4.1 simultaneously to all $\nu_{j}$’s satisfying the hypotheses (not just to a countable family of such $\nu_{j}$’s). So we will not be able to do this by taking intersections of almost sure events naïvely. Rather, we proceed by arguing that the conclusion of the lemma is a property of the type of $(u_{1}(\omega),u_{2}(\omega),\dots)$. We already know that the type of the Haar unitaries converges almost surely to $\mu_{\operatorname{Haar}}$ (there we only had to test countably many formulas). We will show that for each $\nu_{1}$, …, $\nu_{k}$ as in Lemma 4.1, the conclusion of the lemma can be expressed in terms of formulas in $(u_{1},u_{2},\dots)$, and so it depends only the type. Therefore, since the type $\mu_{\operatorname{Haar}}$ can be described only by testing countably many formulas, one can indeed obtain the conclusion of Lemma 4.1 for uncountably many choices of $(\nu_{1},\dots,\nu_{k})$.
The next lemma follows from Lemma 4.1 from purely model-theoretic considerations, specifically the countable saturation of the matrix ultraproduct ${\mathcal{Q}}$. Compare for instance [5, proof of Corollary 9.10]. In the following, a modulus of continuity refers to a continuous increasing function $w:[0,\infty)\to[0,\infty)$ such that $w(0)=0$.
Lemma 4.2.
Let $\mu_{\operatorname{Haar}}\in\mathbb{S}_{(1,1,\dots)}(\operatorname{Th}({\mathcal{Q}}))$ be the almost sure limit of the type of Haar random unitary matrices as $n\to{\mathcal{U}}$. Let $k\in{\mathbb{N}}$, and let $i_{1}\neq i_{2}\neq\dots\neq i_{k}$. For each $j=1$, …, $k$, let $\nu_{j}\in\mathbb{S}_{(1,1)}(\mathrm{T}_{\operatorname{tr}})$ be the type of some pair $(u,x)$ where $u$ is a Haar unitary and $x$ has trace zero, such that $\operatorname{Ent}^{{\mathcal{U}}}(\nu_{j})=0$. Then there exist $2$-variable nonnegative definable predicates $\phi_{1}$, …, $\phi_{k}$ with $\nu_{j}[\phi_{j}]=0$ for each $j$ and there exists a modulus of continuity $w:[0,\infty)\to[0,\infty)$, such that $\mu_{\operatorname{Haar}}$ annihilates the formula $\psi$ given by
(4.1) $\psi^{{\mathcal{Q}}}(u_{1},u_{2},\dots):=\sup_{x_{1},\dots,x_{k}\in B_{1}^{{\mathcal{Q}}}}\left(|\operatorname{tr}^{{\mathcal{Q}}}(x_{1}\dots x_{k})|\dot{-}w(\max_{j}\phi_{j}^{{\mathcal{Q}}}(u_{i_{j}},x_{j}))\right),$
where $a\dot{-}b=\max(a-b,0)$.
Proof.
From Lemma 4.1, we know that if $(u_{1},u_{2},\dots)$ realizes the type $\mu_{\operatorname{Haar}}$ and if $\operatorname{tp}^{{\mathcal{Q}}}(u_{i_{j}},x_{j})=\nu_{j}$ for $j=1$, …, $k$, then $\operatorname{tr}^{{\mathcal{Q}}}(x_{1}\dots x_{k})=0$.
By Lemma 3.2, fix a $2$-variable definable predicate $\phi_{j}\geq 0$ such that for $u,x\in B_{1}^{{\mathcal{Q}}}$, we have $\phi_{j}^{{\mathcal{Q}}}(u,x)=0$ if and only if $\operatorname{tp}^{{\mathcal{Q}}}(u,x)=\nu_{j}$. We claim that for every $\epsilon>0$, there exists a $\delta>0$ such that if $\max_{j}\phi_{j}^{{\mathcal{Q}}}(u_{i_{j}},x_{j})<\delta$, then $|\operatorname{tr}^{{\mathcal{Q}}}(x_{1}\dots x_{k})|<\epsilon$. Suppose for contradiction that this fails. Then there exists some $\epsilon>0$ such that for all $\delta>0$, there exist $x_{1},\dots,x_{k}$ with $\phi_{j}^{{\mathcal{Q}}}(u_{i_{j}},x_{j})<\delta$ for each $j$ but $|\operatorname{tr}^{{\mathcal{Q}}}(x_{1}\dots x_{k})|\geq\epsilon$. Consider the definable predicates $\{\phi_{j}(u_{i_{j}},x_{j})\}_{j=1}^{k}\cup\{\epsilon\dot{-}|\operatorname{tr}(x_{1}\dots x_{k})|\}$ in variables $(x_{1},\dots,x_{k})$ and constants $(u_{1},u_{2},\dots)$. Then this set of formulas is approximately satisfiable, and so by countable saturation of ${\mathcal{Q}}$ (Lemma 3.7), it is satisfiable. That is, there exist $(x_{1},\dots,x_{k})$ with $\phi_{j}(u_{i_{j}},x_{j})=0$ and $|\operatorname{tr}^{{\mathcal{Q}}}(x_{1}\dots x_{k})|\geq\epsilon$. This contradicts the conclusion of Lemma 4.1 that we stated at the beginning of the proof.
Knowing that for every $\epsilon>0$, there exists a $\delta>0$ such that if $\max_{j}\phi_{j}^{{\mathcal{Q}}}(u_{i_{j}},x_{j})<\delta$, then $|\operatorname{tr}^{{\mathcal{Q}}}(x_{1}\dots x_{k})|<\epsilon$, one can construct a continuous increasing function $w:[0,\infty)\to[0,\infty)$ with $w(0)=0$ such that $|\operatorname{tr}^{{\mathcal{Q}}}(x_{1}\dots x_{k})|\leq w(\max_{j}\phi_{j}^{{\mathcal{Q}}}(u_{i_{j}},x_{j}))$ for all $x_{1},\dots,x_{k}\in B_{1}^{{\mathcal{Q}}}$. ∎
Remark 4.3.
Lemma 4.2 is where we rely the fact on working with the full types rather than the existential types. Since the space of existential types is not Hausdorff, one cannot use continuous functions to separate points in that space. Hence, it is necessary to reason about a larger space which is compact and Hausdorff.
Theorem 4.4.
Let $(u_{1},u_{2},\dots)$ be a tuple in ${\mathcal{Q}}$ with type equal to $\mu_{\operatorname{Haar}}$ of Lemma 3.4. For each $i\in{\mathbb{N}}$, let ${\mathcal{A}}_{i}\ni u_{i}$ with $\operatorname{Ent}^{{\mathcal{U}}}({\mathcal{A}}_{i}:{\mathcal{M}})=0$. Then ${\mathcal{A}}_{1}$, ${\mathcal{A}}_{2}$, …are freely independent.
Proof.
Let $i_{1}\neq i_{2}\neq\dots\neq i_{k}$. Let $x_{j}\in{\mathcal{A}}_{i_{j}}$ with $\operatorname{tr}^{{\mathcal{Q}}}(x_{j})=0$, and we will show that $\operatorname{tr}^{{\mathcal{Q}}}(x_{1}\dots x_{k})=0$. Let $\nu_{j}$ be the type of $(u_{j},x_{j})$. Note that by (3.1),
$\operatorname{Ent}^{{\mathcal{U}}}(\nu_{j})=\operatorname{Ent}^{{\mathcal{U}}}(\mathrm{W}^{*}(u_{j},x_{j}):{\mathcal{M}})\leq\operatorname{Ent}^{{\mathcal{U}}}({\mathcal{A}}_{j}:{\mathcal{M}})=0.$
Therefore, by Lemma 4.2, there exists some modulus of continuity $w$ such that $\mu_{\operatorname{Haar}}$ vanishes on the formula $\psi$ of (4.1). In particular, we have
$|\operatorname{tr}^{{\mathcal{Q}}}(x_{1}\dots x_{k})|\dot{-}w(\max_{j}\phi_{j}^{{\mathcal{Q}}}(u_{i_{j}},x_{j}))=0.$
Since $\phi_{j}^{{\mathcal{Q}}}(u_{i_{j}},x_{j})=\nu_{j}[\phi_{j}]=0$, we obtain that $\operatorname{tr}^{{\mathcal{Q}}}(x_{1}\dots x_{k})=0$ as desired. Since this holds for every alternating string $i_{1}\neq\dots\neq i_{k}$ and all trace-zero $x_{j}\in{\mathcal{A}}_{i_{j}}$, we have that ${\mathcal{A}}_{1}$, ${\mathcal{A}}_{2}$, …are freely independent as desired. ∎
Proof of Theorem C.
Fix an outcome $\omega$ in the almost sure event where the type $(U_{1}^{(n)},U_{2}^{(n)},\dots)$ converges to $\mu_{\operatorname{Haar}}$, or in other words the type of $(u_{1}(\omega),u_{2}(\omega),\dots)$ is $\mu_{\operatorname{Haar}}$. Let ${\mathcal{P}}_{j}$ be the Pinsker algebra containing $u_{j}$. Since $\operatorname{Ent}^{{\mathcal{U}}}({\mathcal{P}}_{j}:{\mathcal{Q}})\leq h({\mathcal{P}}_{j}:{\mathcal{Q}})=0$ by (3.2), the ${\mathcal{P}}_{j}$’s are freely independent by Theorem 4.4. ∎
Remark 4.5 (On the theory of matrix ultraproducts).
Fix $2$-variable types $\nu_{j}$ with $\operatorname{Ent}^{{\mathcal{U}}}(\nu_{j})=0$, and let $\psi$ be as in Lemma 4.2. Note that $\psi$ only depends on the finitely many of $u_{i}$’s, and since the unitaries form a definable set, the following is a definable predicate with no free variables:
$\gamma=\inf_{u_{1},u_{2},\dots\text{ unitary}}\psi(u_{1},u_{2},\dots).$
Then the matrix ultraproduct ${\mathcal{Q}}$ satisfies the sentence $\gamma^{{\mathcal{Q}}}=0$ since $\psi^{{\mathcal{Q}}}$ vanishes on the unitaries with type $\mu_{\operatorname{Haar}}$. By similar reasoning as in Lemma 4.2, the free independence of the commutants of the $u_{j}$’s from Theorem B leads to the following statement: Fix a word $i_{1}\neq i_{2}\neq\dots\neq i_{k}$. There is some modulus of continuity $w$ such that
$\inf_{u_{1},u_{2},\dots\in{\mathcal{Q}}\text{ unitary}}\sup_{x_{1},\dots,x_{k}\in B_{1}^{{\mathcal{Q}}}}\left(|\operatorname{tr}^{{\mathcal{Q}}}(x_{1}\dots x_{k})\dot{-}w(\max_{j}\lVert[u_{i_{j}},x_{j}]\rVert_{2})\right)=0.$
[…]
4.2. Free independence phenomena in ultraproducts of free products
Proof of Theorem F.
First, to prove claims (1), (2), (3), it suffices to consider the case where $I$ is countable. Indeed, for (1), we have to prove that for every $i_{1}\neq i_{2}\neq\dots\neq i_{k}$ if $x_{j}\in{\mathcal{A}}_{i_{j}}$ with trace zero, then $\operatorname{tr}^{{\mathcal{M}}^{{\mathcal{V}}}}(x_{1}\dots x_{k})=0$. Let $I_{0}$ be a countable subset of $I$ containing the indices $i_{1}$, …, $i_{k}$. Then view ${\mathcal{M}}$ as a free product of countably many algebras ${\mathcal{M}}_{0}*(*_{i\in I_{0}}{\mathcal{M}}_{i})$, where ${\mathcal{M}}_{0}=*_{i\in I\setminus I_{0}}{\mathcal{M}}_{i}$. By applying the conclusions of the theorem for this free product decomposition, we obtain $\operatorname{tr}^{{\mathcal{M}}^{{\mathcal{V}}}}(x_{1}\dots x_{k})=0$. Hence, in the remainder of the proof assume without loss of generality that $I={\mathbb{N}}$.
Now because ${\mathcal{M}}_{i}$ is assumed to be diffuse and Connes embeddable, there exists an ultrafilter ${\mathcal{V}}^{\prime}$ on some infinite index set [^3] Since ${\mathcal{M}}_{i}$ is not necessarily separable, the index set for ${\mathcal{V}}^{\prime}$ may be uncountable. such that ${\mathcal{M}}_{i}$ embeds into ${\mathcal{R}}^{{\mathcal{V}}^{\prime}}$ for each $i\in I$, where ${\mathcal{R}}$ is the unique separable hyperfinite II1 factor. Since we assumed that ${\mathcal{A}}_{i}\cap{\mathcal{M}}_{i}$ is diffuse, let $a_{i}$ be a Haar unitary in ${\mathcal{M}}_{i}$. We may assume without loss of generality that inclusion ${\mathcal{M}}_{i}\to{\mathcal{R}}^{{\mathcal{V}}^{\prime}}$ sends $a_{i}$ into the diagonal copy of ${\mathcal{R}}$ in ${\mathcal{R}}^{{\mathcal{V}}^{\prime}}$; this is because all Haar unitaries in ${\mathcal{R}}^{{\mathcal{V}}^{\prime}}$ are conjugate to each other.
Let ${\mathcal{S}}=*_{i\in I}{\mathcal{R}}\cong L(F_{\infty})$. For convenience of notation, we will denote by ${\mathcal{R}}_{i}$ the $i$th copy of ${\mathcal{R}}$ in ${\mathcal{S}}$. The inclusions ${\mathcal{R}}_{i}\to{\mathcal{S}}$ yields an inclusion ${\mathcal{R}}_{i}^{{\mathcal{V}}^{\prime}}$ into ${\mathcal{S}}^{{\mathcal{V}}^{\prime}}$, and the ${\mathcal{R}}_{i}^{{\mathcal{V}}^{\prime}}$’s are freely independent in ${\mathcal{S}}^{{\mathcal{V}}^{\prime}}$ by a straightforward limiting argument. Hence, we have mappings
$*_{i\in I}{\mathcal{M}}_{i}\to*_{i\in I}{\mathcal{R}}_{i}^{{\mathcal{V}}^{\prime}}\to(*_{i\in I}{\mathcal{R}}_{i})^{{\mathcal{V}}^{\prime}}.$
Thus also we have inclusions
${\mathcal{M}}^{{\mathcal{V}}}=(*_{i\in I}{\mathcal{M}}_{i})^{{\mathcal{V}}}\to(*_{i\in I}{\mathcal{R}}_{i}^{{\mathcal{V}}^{\prime}})^{{\mathcal{V}}}\to((*_{i\in I}{\mathcal{R}}_{i})^{{\mathcal{V}}^{\prime}})^{{\mathcal{V}}}\cong(*_{i\in I}{\mathcal{R}}_{i})^{{\mathcal{W}}}={\mathcal{S}}^{{\mathcal{W}}},$
where ${\mathcal{W}}$ is a certain ultrafilter on the product of the index sets for ${\mathcal{V}}^{\prime}$ and ${\mathcal{V}}$. We have also arranged that the Haar unitary $a_{i}$ in ${\mathcal{M}}_{i}\cap{\mathcal{A}}_{i}\subseteq(*_{i\in I}{\mathcal{M}}_{i})^{{\mathcal{V}}}$ is contained in ${\mathcal{R}}_{i}\subseteq{\mathcal{R}}_{i}^{{\mathcal{W}}}\subseteq{\mathcal{S}}^{{\mathcal{W}}}$.
Note that $h({\mathcal{M}}_{i}:{\mathcal{S}}^{{\mathcal{W}}})\leq h({\mathcal{R}}_{i}^{{\mathcal{V}}^{\prime}}:{\mathcal{S}}^{{\mathcal{W}}})=0$ since ${\mathcal{R}}^{{\mathcal{V}}^{\prime}}$ has property Gamma. Let ${\mathcal{P}}_{i}$ be the Pinsker algebra of ${\mathcal{M}}_{i}$ in ${\mathcal{S}}^{{\mathcal{W}}}$, that is, the unique maximal subalgebra with $h({\mathcal{P}}_{i}:{\mathcal{S}}^{{\mathcal{W}}})=0$. Of course, we have that ${\mathcal{R}}_{i}\subseteq{\mathcal{P}}_{i}$ since ${\mathcal{R}}_{i}\cap{\mathcal{M}}_{i}$ is diffuse.
We claim that the ${\mathcal{P}}_{i}$’s are freely independent (the three claims in the theorem statement will follow from this, as we explain at the end of the proof). It suffices to show free independence of any separable subalgebras ${\mathcal{B}}_{i}$ inside ${\mathcal{P}}_{i}$ such that ${\mathcal{R}}_{i}\subseteq{\mathcal{B}}_{i}$. Fix a free ultrafilter ${\mathcal{U}}$ on ${\mathbb{N}}$ and let ${\mathcal{Q}}=\prod_{n\to{\mathcal{U}}}M_{n}({\mathbb{C}})$. Let $A^{(n)}=\operatorname{diag}(1,\zeta_{n},\dots,\zeta_{n}^{n-1})$, where $\zeta_{n}=e^{2\pi i/n}$, and let $a=[A^{(n)}]_{n\in{\mathbb{N}}}\in{\mathcal{Q}}$. Fix an embedding $\pi_{i}:{\mathcal{R}}_{i}\to{\mathcal{Q}}$. Since all Haar unitaries in ${\mathcal{Q}}$ are conjugate, assume without loss of generality that $\pi_{i}(a_{i})=a$. Let $(V_{i}^{(n)})_{i\in{\mathbb{N}}}$ be independent $n\times n$ Haar random unitaries. By Proposition 2.2 and Lemma 3.4, we have that almost surely $\operatorname{tp}^{M_{n}({\mathbb{C}})}(V_{1}^{(n)}A^{(n)}(V_{1}^{(n)})^{*},V_{2}^{(n)}A^{(n)}(V_{2}^{(n)})^{*},\dots)$ converges to the type $\mu_{\operatorname{Haar}}$ of Lemma 3.4. Thus, fix an outcome $\omega$ where this occurs, let $v_{i}=[V_{i}^{(n)}(\omega)]_{n\in{\mathbb{N}}}\in{\mathcal{Q}}$, and let $u_{i}=v_{i}av_{i}^{*}$, so that the type of $(u_{1},u_{2},\dots)$ is $\mu_{\operatorname{Haar}}$. Note that $v_{i}\pi_{i}({\mathcal{R}}_{i})v_{i}^{*}$ contains $u_{i}$ and is amenable; therefore, using (3.2) and [29, §2.3, Property 4].
$0\leq\operatorname{Ent}^{{\mathcal{U}}}(v_{i}\pi_{i}({\mathcal{R}}_{i})v_{i}^{*}:{\mathcal{Q}})\leq h^{{\mathcal{U}}}(v_{i}\pi_{i}({\mathcal{R}}_{i}))v_{i}^{*}:{\mathcal{Q}})\leq h(v_{i}\pi_{i}({\mathcal{R}}_{i}))v_{i}^{*}:{\mathcal{Q}})\leq 0$
Thus, by Theorem 4.4, the algebras $v_{i}\pi_{i}({\mathcal{R}}_{i})v_{i}^{*}$ are freely independent. In particular, there is an embedding $\pi:{\mathcal{S}}=*_{i\in I}{\mathcal{R}}_{i}\to{\mathcal{Q}}$ such that $\pi|_{{\mathcal{R}}_{i}}=\operatorname{ad}_{v_{j}}\circ\pi_{i}$.
By the downward Löwenheim-Skolem theorem (Proposition 3.6), choose some separable elementary substructure $\widehat{{\mathcal{S}}}$ of ${\mathcal{S}}^{{\mathcal{W}}}$ containing all the ${\mathcal{B}}_{i}$’s. We claim that $\pi$ extends to an embedding $\widehat{\pi}:\widehat{{\mathcal{S}}}\to{\mathcal{Q}}$. In more detail, let $\mathbf{x}$ and $\mathbf{y}$ be countable tuple generating ${\mathcal{S}}$ and $\widehat{{\mathcal{S}}}$ respectively, and assume they are elements in the unit ball. Since ${\mathcal{S}}\subseteq\widehat{{\mathcal{S}}}\subseteq{\mathcal{S}}^{{\mathcal{W}}}$, then for every $\epsilon>0$ and for any finitely many polynomials $p_{1}$, …, $p_{\ell}$ of countably many variables, there exists some tuple $\mathbf{z}$ from the unit ball in ${\mathcal{S}}$ such that
$|\operatorname{tr}^{{\mathcal{S}}}(p_{j}(\mathbf{x},\mathbf{z}))-\operatorname{tr}^{{\mathcal{S}}^{{\mathcal{W}}}}(p_{j}(\mathbf{x},\mathbf{y}))|<\epsilon.$
In particular, there exists $\mathbf{w}$ in ${\mathcal{Q}}$ (namely $\mathbf{w}=\pi(\mathbf{z})$) such that
$|\operatorname{tr}^{{\mathcal{Q}}}(p_{j}(\pi(\mathbf{x}),\mathbf{w}))-\operatorname{tr}^{{\mathcal{S}}^{{\mathcal{W}}}}(p_{j}(\mathbf{x},\mathbf{y}))|<\epsilon.$
Hence, by countable saturation of ${\mathcal{Q}}$ (Lemma 3.7), there exists $\mathbf{w}$ in ${\mathcal{Q}}$ such that
$\operatorname{tr}^{{\mathcal{Q}}}(p(\pi(\mathbf{x}),\mathbf{w}))=\operatorname{tr}^{{\mathcal{M}}^{{\mathcal{V}}}}(p(\mathbf{x},\mathbf{y}))$
for all non-commutative polynomials $p$. Hence, there is a trace-preserving embedding $\widehat{\pi}:\widehat{{\mathcal{S}}}\to{\mathcal{Q}}$ such that $\widehat{\pi}(\mathbf{x})=\pi(\mathbf{x})$ and $\widehat{\pi}(\mathbf{y})=\mathbf{w}$.
Now
$h(\widehat{\pi}({\mathcal{B}}_{i}):{\mathcal{Q}})\leq h(\widehat{\pi}({\mathcal{B}}_{i}):\widehat{\pi}(\widehat{{\mathcal{S}}}))=h({\mathcal{B}}_{i}:\widehat{{\mathcal{S}}})=h({\mathcal{B}}_{i}:{\mathcal{S}}^{{\mathcal{W}}})\leq h({\mathcal{P}}_{i}:{\mathcal{S}}^{{\mathcal{W}}})=0;$
here the equality $h({\mathcal{B}}_{i}:\widehat{{\mathcal{S}}})=h({\mathcal{B}}_{i}:{\mathcal{S}}^{{\mathcal{W}}})$ follows because $h({\mathcal{B}}_{i}:\widehat{{\mathcal{S}}})$ only depends on the existential type of generators of ${\mathcal{B}}_{i}$ in $\widehat{{\mathcal{S}}}$, and this is the same as its existential type in ${\mathcal{S}}^{{\mathcal{W}}}$ since $\widehat{{\mathcal{S}}}$ is an elementary substructure. Since $\widehat{\pi}({\mathcal{B}}_{i})$ by construction contains $\operatorname{ad}_{v_{i}}\circ\pi_{i}({\mathcal{R}}_{i})$ and hence contains $v_{i}av_{i}^{*}=u_{i}$, we can apply Theorem 4.4 to obtain that the algebras $\pi_{i}({\mathcal{B}}_{i})$ are freely independent in ${\mathcal{Q}}$. This means also that the ${\mathcal{B}}_{i}$’s are freely independent in ${\mathcal{S}}^{{\mathcal{W}}}$.
Therefore, we have shown that the Pinsker algebras ${\mathcal{P}}_{i}$ are freely independent. The claims of the theorem now follow quickly from this, together with the properties of $h$:
  1. (1)
    If ${\mathcal{A}}_{i}\subseteq{\mathcal{M}}^{{\mathcal{V}}}\subseteq{\mathcal{S}}^{{\mathcal{W}}}$ and ${\mathcal{A}}_{i}\cap{\mathcal{M}}_{i}$ is diffuse and $h({\mathcal{A}}_{i}:{\mathcal{M}}^{{\mathcal{V}}})=0$, then we obtain using [29, §2.3, Property 3] that $h({\mathcal{A}}_{i}:{\mathcal{S}}^{{\mathcal{W}}})\leq h({\mathcal{A}}_{i}:{\mathcal{M}}^{{\mathcal{V}}})=0$, and therefore ${\mathcal{A}}_{i}$ is contained inside the Pinsker algebra ${\mathcal{P}}_{i}$. Thus, the ${\mathcal{A}}_{i}$’s are freely independent.
  2. (2)
    Similar to point (1), it suffices to note that ${\mathcal{C}}_{i}$ is contained in the Pinsker algebra ${\mathcal{P}}_{i}$, which follows from [38, Fact 2.9] as noted in the introduction at Corollary E.
  3. (3)
    Note that the von Neumann ${\mathcal{N}}_{i}$ generated by the wq-normalizer of ${\mathcal{M}}_{i}$ in ${\mathcal{M}}$ is contained in the von Neumann algebra $\tilde{{\mathcal{N}}}_{i}$ generated by wq-normalizer of ${\mathcal{M}}_{i}$ in ${\mathcal{S}}^{{\mathcal{W}}}$. By [29, §2.3, Properties 3 and 9], we have $h(\tilde{{\mathcal{N}}}_{i}:{\mathcal{S}}^{{\mathcal{W}}})=h({\mathcal{M}}_{i}:{\mathcal{S}}^{{\mathcal{W}}})\leq h({\mathcal{R}}_{i}^{{\mathcal{V}}^{\prime}}:{\mathcal{S}}^{{\mathcal{W}}})=0$. Since $\tilde{{\mathcal{N}}}_{i}\cap{\mathcal{M}}_{i}$ is diffuse, we have that $\tilde{{\mathcal{N}}}_{i}$ is contained in the Pinsker algebra ${\mathcal{P}}_{i}$. Thus, the ${\mathcal{N}}_{i}$’s are freely independent as desired. ∎
Declarations
Data availability statement: This manuscript has no associated data.
Funding and/or Conflicts of interests/Competing interests: DJ acknowledges funding from the grant ‘‘Logic and $\mathrm{C}^{*}$-algebras’’ from National Sciences and Engingeering Research Council of Canada; the Independent Research Fund of Denmark, grant 1026-00371B; and the Horizon Europe Marie Skłodowska-Curie Action FREEINFOGEOM. SKE was funded by the National Science Foundation (US) grant DMS 2350049. We thank the American Institute of Mathematics and the NSF for funding DJ’s visit to California in April 2024 where this paper first was developed.
The authors have no relevant financial or non-financial interests to disclose.